### v1

# Factoriality and birational rigidity of two families of singular quartic three-folds

Factoriality and birational rigidity
of two families of singular quartic three-folds
A.V.Pukhlikov
In this paper we study two families of three-dimensional quartics in the complex projective space ${\mathbb{P}}^{4}$: hypersurfaces with a unique quadratic singularity of rank 3, which is resolved by two blowups, and hypersurfaces with two quadratic singularities of rank 3 and 4. Both families have codimension 3 in the natural parameter space. For a Zariski general quartic in each of these families we prove factoriality and birational rigidity and describe its group of birational self-maps.
Bibliography: 15 items.
AMS classification: 14E05, 14E07
Key words: Fano variety, birational rigidity, birational automorphism, maximal singularity, linear system, multiplicity.
Introduction
0.1. The main result. The present paper is a follow up of [1]. We prove factoriality and birational rigidity and describe the group of birational automorphisms of two classes of three-dimensional quartics in ${\mathbb{P}}^{4}$ (the ground field is the field of complex numbers $\mathbb{C}$}:
(1) the quartics $V\subset{\mathbb{P}}^{4}$ with a unique singular point $o_{1}\in V$, which is a quadratic singularity of rank 3, the blowup $\varphi_{1}\colon V_{1}\to V$ of which yields a variety $V_{1}$ with a unique singularity, a non-degenerate quadratic singularity $o_{2}\in V_{1}$ (so that the blowup $\varphi_{2}\colon V_{2}\to V_{1}$ of the point $o_{2}$ gives a non-singular variety $V_{2}$), — in this paper we will call such quartics the quartics of the first type,
(2) the quartics $V\subset{\mathbb{P}}^{4}$ with two quadratic singular points $o_{1}$, $o_{2}$, where the rank of the point $o_{1}$ is 3, and the rank of the point $o_{2}$ is 4, and the blowup of the two points $o_{1}$, $o_{2}$ gives a non-singular variety, we will call such hypersurfaces the quartics of the second type.
In both cases we will assume that the quartic $V$ satisfies certain conditions of general position, see Subsection 0.2. Both families have codimension 3 in the (projectively) 69-dimensional space of three-dimensional quartics. Thus the present paper can be considered as the next step (after [1]) in the proof of the old conjecture [2, 3] that factorial three-dimensional quartics with isolated singularities not worse than $cA_{1}$ are birationally rigid; or as the next step in the description of the birational geography of three-dimensional quartics, see Subsection 0.5 below.
0.2. Quadratic singularities. Our notations are as close as possible to the notations of [1]. Let ${\mathbb{A}}^{4}\subset{\mathbb{P}}^{4}$ be an affine chart with the coordinates $(z_{1},z_{2},z_{3},z_{4})$ and
$f(z_{*})=q_{2}(z_{*})+q_{3}(z_{*})+q_{4}(z_{*})$
a polynomial of degree 4, where $q_{i}(z_{*})$ are homogeneous of degree $i\in\{2,3,4\}$, where $q_{2}=z_{1}^{2}+z_{2}^{2}+z_{3}^{2}$ and the set of common zeros of the polynomials $q_{2}$, $q_{3}$, $q_{4}$ on ${\mathbb{P}}^{3}_{(z_{1}:z_{2}:z_{3}:z_{4})}$ consists of 24 distinct points. Let $\pi_{1}\colon X_{1}\to{\mathbb{P}}^{4}$ be the blow up of the point $o_{1}=(0,0,0,0)$, denote by the symbol ${\mathbb{E}}_{1}$ its exceptional divisor, ${\mathbb{E}}_{1}\cong{\mathbb{P}}^{3}_{(z_{1}:z_{2}:z_{3}:z_{4})}$. The symbol $V$ stands for the hypersurface $\{f=0\}\subset{\mathbb{P}}^{4}$, and the symbol $V_{1}$ — for its strict transform on $X_{1}$, so that $\varphi_{1}\colon V_{1}\to V$ is the blow-up of the point $o_{1}$ (now on $V$). Let $Q_{1}=V_{1}\cap{\mathbb{E}}_{1}$ be the exceptional quadric. The only candidate for a singular point of the variety $V_{1}$ on $Q_{1}$ is the vertex of the quadratic cone $Q_{1}$, i.e. the point
$o_{2}=(0:0:0:1).$
We define the quartic of the first type, assuming that $o_{2}\in V_{1}$ is a non-degenerate (nodal) quadratic point: in terms of the polynomial $f$ it means that
$q_{3}(0,0,0,1)=0,$
however,
$4q_{4}(0,0,0,1)-\sum_{i=1}^{3}\left(\frac{\partial^{3}q_{3}}{\partial z_{i}\partial z_{4}^{2}}(0,0,0,1)\right)^{2}\neq 0$
(the equality of the left hand side to zero means precisely that the point $o_{2}$ is a quadratic singularity of rank 3). Therefore, the singular point $o_{1}\in V$ is resolved by two blow-ups: let $\pi_{2}\colon X_{2}\to X_{1}$ be the blow-up of the point $o_{2}$ with the exceptional divisor ${\mathbb{E}}_{2}\cong{\mathbb{P}}^{3}$ and $V_{2}\subset X_{2}$ the strict transform of the hypersurface $V_{1}$, that is, $\varphi_{2}\colon V_{2}\to V_{1}$ is the blow-up of the point $o_{2}$ on $V_{1}$, then $V_{2}\cap{\mathbb{E}}_{2}=Q_{2}$ is a non-degenerate quadric. We assume that
$q_{4}(0,0,0,1)\neq 0$
(geometrically it means that the line $o_{1},o_{2}]\subset{\mathbb{P}}^{4}$, passing through the point $o_{1}$ in the direction $o_{2}$, is not contained in $V$). We assume that $\{o_{1}\}=\mathop{\rm Sing}V$. In order to reduce the number of cases to be considered, we will assume that the quartic $V$ satisfies some additional conditions of general position, which will be stated in §§3, 5.
The quartics of the second type are defined by the conditions $q_{3}(0,0,0,1)\neq 0$ (so that $V_{1}$ is non-singular at the points of the quadric $Q_{1}$), however, $V$ has a non-degenerate quadratic singularity $o_{2}\neq o_{1}$, and moreover,
$\mathop{\rm Sing}V=\{o_{1},o_{2}\}.$
In addition we assume that the line $o_{1},o_{2}]\subset{\mathbb{P}}^{4}$, containing these points, is not contained in $V$. Apart from that, certain conditions of general position, simplifying our arguments, are stated in §§2,4.
In order to make our notations uniform, in the case of a quartic of the second type let $\pi_{2}\colon X_{2}\to X_{1}$ be the blow-up of the point $o_{2}$ (which now is not infinitely near to $o_{1}$), ${\mathbb{E}}_{2}\cong{\mathbb{P}}^{3}$ the exceptional divisor, $V_{2}\subset X_{2}$ the strict transform and $Q_{2}=V_{2}\cap{\mathbb{E}}_{2}$ a non-degenerate quadric. The second blow-up $\pi_{2}$ does not affect the cone $Q_{1}$ (as $o_{2}\not\in Q_{1}$) and one may write $Q_{1}\subset X_{2}$; obviously, $Q_{1}\cap Q_{2}=\emptyset$.
Remark 0.1. For quartics of the first type the condition $o_{1},o_{2}]\not\subset V$ follows from the assumption that on $V$ there are precisely 24 distinct lines, passing through the point $o_{1}$: if $q_{4}(0,0,0,1)=0$, then the vertex of the cone $Q_{1}$ (the point $o_{2}$) comes into the cycle of the scheme-theoretic intersection of the surfaces
$\{q_{2}=0\},\quad\{q_{3}=0\},\quad\{q_{4}=0\}$
in ${\mathbb{E}}_{1}\cong{\mathbb{P}}_{(z_{1}:z_{2}:z_{3}:z_{4})}$ with multiplicity $\geqslant 2$, so that on $V$ there are at most 23 distinct lines, passing through the point $o_{1}$. For quartics of the second type the condition $o_{1},o_{2}]\not\subset V$ is assumed to hold in order to reduce the number of cases to inspect in the proof of the plane section lemma, see §4.
0.3. Factoriality and birational rigidity. We have the equality
$\mathop{\rm Pic}X_{2}={\mathbb{Z}}H_{{\mathbb{P}}}\oplus{\mathbb{Z}}{\mathbb{E}}_{1}\oplus{\mathbb{Z}}{\mathbb{E}}_{2}$
(which holds for quartics of both types), where we omit the symbols of pull back $\pi_{1}^{*}$, $\pi_{2}^{*}$, and $H_{{\mathbb{P}}}$ is the class of a hyperplane in ${\mathbb{P}}^{4}$. Obviously,
${\mathbb{Z}}H\oplus{\mathbb{Z}}Q_{1}\oplus{\mathbb{Z}}Q_{2}\subset\mathop{\rm Pic}V_{2}$
(here $H$ is the class of a hyperplane section of $V$).
Theorem 0.1. The following equality holds:
$\mathop{\rm Pic}V_{2}={\mathbb{Z}}H\oplus{\mathbb{Z}}Q_{1}\oplus{\mathbb{Z}}Q_{2},$
so that $V\subset{\mathbb{P}}^{4}$ is a factorial hypersurface with $\mathop{\rm Cl}V=\mathop{\rm Pic}V={\mathbb{Z}}H$.
Let $\Sigma$ be a mobile linear system on $V$ (that is, $\mathop{\rm codim}\mathop{\rm Bs}\Sigma\geqslant 2$, the system $\Sigma$ has no fixed components). By Theorem 0.1, there is $n[\Sigma]\geqslant 1$, such that $\Sigma\subset|n[\Sigma]\cdot H|$. We say that $\Sigma$ has no maximal singularities, if the pair $\left(V,\frac{1}{n[\Sigma]}D\right)$ is canonical, that is, for every excetional divisor $E$ over $V$ the inequality
$\mathop{\rm ord}\nolimits_{E}D\leqslant n[\Sigma]\cdot a(E)$
holds, where $D\in\Sigma$ is a general divisor and $a(E)$ is the discrepancy.
Theorem 0.2. The quartic $V$ is a birationally rigid variety: for every mobile linear system $\Sigma$ there is a birational automorphism $\chi\in\mathop{\rm Bir}V$, such that the linear system $(\chi^{-1})_{*}\Sigma$ has no maximal singularities. On $V$ there are no structures of a conic bundle and fibrations into del Pezzo surfaces, and every birational map $V\dashrightarrow V^{\prime}$ onto a three-dimensional Fano variety with $\mathbb{Q}$}-factorial terminal singularities is an isomorphism.
(It is well known that the second claim of Theorem 0.2 follows from the first one.)
The third main result of the present paper is an explicit description of the group of birational automorphisms $\mathop{\rm Bir}V$. In contrast to the first two main results, its statement is not the same for the quartics of the first and second type. This is because the second singularity of a quartiv of the first type is an infinitely near point and does not generate any additional birational automorphisms.
For a quartic $V$ of the first type let $\tau_{0}$ be the Galois involution, generated by the projection $V\dashrightarrow{\mathbb{P}}^{3}$ from the point $o_{1}$ and $\tau_{1}$, …, $\tau_{24}$ are the birational involutions, linked to the 24 lines on $V$, passing through the point $o_{1}$, see [4, 1]. The subgroup in $\mathop{\rm Bir}V$, generated by these 25 involutions, will be denoted by the symbol $B(V)$. For a quartic $V$ of the second type let $\tau_{1,0},\dots,\tau_{1,24}$ have the same meaning as the 25 involutions, listed above, $\tau_{2,0}$ the Galois involution, generated by the projection from the point $o_{2}$, and $\tau_{2,i}$, where $i\in\{1,\dots,L(o_{2})\}$, are the birational involutions, linked to the lines on $V$, passing through the point $o_{2}$. We know that $L(o_{2})\leqslant 24$.
Proposition 0.1. For a general quartic of the second type the equality $L(o_{2})\leqslant 24$ holds.
We denote the subgroup in $\mathop{\rm Bir}V$, generated by these $26+L(o_{2})$ involutions, by the symbol $B(V)$.
Theorem 0.3. The subgroup $B(V)$ is normal in $\mathop{\rm Bir}V$, and there is the exact sequence
$1\to B(V)\to\mathop{\rm Bir}V\to\mathop{\rm Aut}V\to 1,$
where the subgroup $\mathop{\rm Aut}V\subset\mathop{\rm Bir}V$ is finite, for a general (in its family) quartic it is trivial. The involutions, generating the subgroup $B(V)$, form a system of free generators:
$B(V)=\mathop{*}\limits_{i=0}^{24}\langle\tau_{i}\rangle$
for quartics of the first type and
$B(V)=\mathop{*}\limits_{i=0}^{24}\langle\tau_{1,i}\rangle*\,\,\mathop{*}\limits_{i=0}^{L(o_{2})}\langle\tau_{2,i}\rangle$
for quartics of the second type.
Remark 0.2. In the semi-direct product for the group $\mathop{\rm Bir}V$ the finite subgroup $\mathop{\rm Aut}V$ acts on $B(V)$ by conjugation, preserving the point $o_{1}$ and permuting the lines, passing through the point $o_{1}$, for quartics of the first type, and preserving both points $o_{1}$, $o_{2}$ and permuting the lines passing through these points (iside each of the two groups of lines). Thus in the first family of quartics the subgroup $B(V)$ is freely generated by 25 involutions and in the second family for a general quartic it is freely generated by 50 involutions.
0.4. The structure of the paper. In §1 we prove that quartics of the first and second type are factorial (Theorem 0.1). In §2 we start the proof of Theorems 0.2 and 0.3: we recall the concept of a maximal singularity of a mobile linear system of divisors on the quartic $V$ and carry out the standard preparatory work (list the possible types of maximal singularities and exclude some of them). The claims of Theorems 0.2 and 0.3 for quartics of the first type are reduced to two facts: the plane section lemma (Proposition 2.2) and the claim that the centre of a maximal singularity on $V_{1}$ can not be a generator of the cone $Q_{1}$ (Proposition 2.4). For quartics of the second type it is sufficient to show the plane section lemma and a few close claims, some of which are shown in §2, and the others are considered in §4.
In §3 we show the plane section lemma for quartics of the first type. We go through all options for quartics of general position (in that family) and a few additional examples. Apart from that, we show some auxiliary claims that are used later when we exclude generators of the cone $Q_{1}$ in §5 (this is the hardest part of the proof of Theorems 0.2 and 0.3).
In §4 we show the plane section lemma for quartics of the second type. Here we also prove Proposition 2.5. Its proof requires the same technique as that of the plane section lemma.
Finally, in §5 we show the key intermediate claim in the proof of the birational rigidity of quartics of the first type: a generator $R$ of the cone $Q_{1}$ can not be the centre of a maximal singularity. We use both the quadratic method (the technique of counting multiplicities for the self-intersection of the system $\Sigma$), and the linear one (similar to the proof of the plane section lemma).
0.5. Birational geography of three-dimensional quartics. Irreducible quartic hypersurfaces in ${\mathbb{P}}^{4}$ are parameterized by points of a Zariski open subset in the projective space ${\mathbb{P}}^{69}$, that is, by points of a quasi-projective 69-dimensional variety. In [1, Subsection 0.4] the problem of studying the ‘‘birational geography’’ of three-dimensional quartics was formulated: to break the natural parameter space into disjoint quasi-projective subvarieties, consisting of quartics with a similar birational type. The natural ‘‘measure of size’’ of these subvarieties is their dimension (or the codimension in ${\mathbb{P}}^{69}$). The seminal paper [5] that started the theory of birational rigidity described the birational type of non-singular quartics, reducing the subsequent work to investigating the 68-dimensional subset of singular quartics, that is, to the subset of codimension 1. It follows from [4] that the birational rigidity still holds in codimension 1, too (although the group of birational automorphisms becomes very large). It follows from [6] (taking into account [7, 8, 9]) and [1] that the codimension of the complement to the set of factorial birationally rigid three-dimensional quartics is at least 3. Non-rigid but ‘‘almost rigid’’ quartics were studies in [2], see also [3]. For the families and types of singularities that are already accessible for the modern techniques (however, often very hard to study), see. [1, Subsection 0.4]. Note also the paper [10], which it is natural to consider as a part of the birational geography of Fano double spaces of dimension 3, branched over a sextic surface in ${\mathbb{P}}^{3}$.
1 Factoriality
In this section we prove Theorem 0.1.
1.1. Preliminary remarks. We use the notations of Subsections 0.2, 0.3, considering in detail quartics of the first type. Note, in the first place, that the proof of factoriality, presented in [2, Remark 1.2], works for a quartics of the first type $V$, too, as by construction the homogeneous polynomial $q_{4}(z_{*})$ is arbitrary: the section of $V$ by the hypersurface ‘‘at infinity’’ is precisely $\{q_{4}=0\}$. Therefore, for a general quartic of the first type its section ‘‘at infinity’’ has Picard number 1. However, we give a different proof of factoriality that does not use this fact.
Note, first of all, that on $X_{2}$ the class of the hypersurface $V_{2}$ is $4H_{\mathbb{P}}-2{\mathbb{E}}_{1}-2{\mathbb{E}}_{2}$. This class is nef and big, but not ample, so that the Lefschetz theorem does not apply (in contrast to [1]). Furthermore, it is easy to see that all cohomology spaces
$H^{1}({\mathbb{P}}^{4},{\cal O}_{{\mathbb{P}}^{4}}),\quad H^{1}(X_{1},{\cal O}_{X_{1}}),\quad H^{1}(X_{2},{\cal O}_{X_{2}}),$
$H^{2}({\mathbb{P}}^{4},{\cal O}_{{\mathbb{P}}^{4}}),\quad H^{2}(X_{1},{\cal O}_{X_{1}}),\quad H^{2}(X_{2},{\cal O}_{X_{2}}),$
and the six similar spaces for $V$, $V_{1}$, $V_{2}$ are trivial, so that (by the explonential sequences for $X_{2}$ and $V_{2}$) there are the natural isomorphisms
$\mathop{\rm Pic}X_{2}\otimes{\mathbb{C}}\cong H^{1}(X_{2},\Omega_{X_{2}})$
and
$\mathop{\rm Pic}V_{2}\otimes{\mathbb{C}}\cong H^{1}(V_{2},\Omega_{V_{2}}),$
and moreover, the restriction of 1-forms onto the non-singular hypersurface $V_{2}$ determines the inclusion
$\rho\colon H^{1}(X_{2},\Omega_{X_{2}})\to H^{1}(V_{2},\Omega_{V_{2}})$
and we have to show that $\rho$ is an isomorphism.
1.2. Exact sequences. Consider the exact sequence
$0\to{\cal N}^{*}_{V_{2}/X_{2}}={\cal O}_{X_{2}}(-V_{2})|_{V_{2}}\to\Omega_{X_{2}}|_{V_{2}}\to\Omega_{V_{2}}\to 0.$
From the obvious equalities $H^{0}(V_{2},\Omega_{V_{2}})=0$ (so that $H^{0}(V_{2},\Omega_{X_{2}}|_{V_{2}})=0$, too) and
$H^{1}(V_{2},{\cal O}_{V_{2}}(-V_{2}))=H^{2}(V_{2},{\cal O}_{V_{2}}(-V_{2}))=0$
(since $4H-2Q_{1}-2Q_{2}$ is a nef and big divisor on $V_{2}$) we get the natural isomorphism
$\gamma\colon H^{1}(V_{2},\Omega_{X_{2}}|_{V_{2}})\to H^{1}(V_{2},\Omega_{V_{2}}).$
Now consider the exact sequence
$0\to{\cal O}_{X_{2}}(-V_{2})\to{\cal O}_{{X_{2}}}\to{\cal O}_{{V_{2}}}\to 0,$
tensored by $\Omega_{X_{2}}$. We get the exact sequence
$0\to H^{1}(X_{2},\Omega_{X_{2}}(-V_{2}))\to H^{1}(X_{2},\Omega_{X_{2}})\to H^{1}(V_{2},\Omega_{X_{2}}|_{V_{2}})\to$
$\to H^{2}(X_{2},\Omega_{X_{2}}(-V_{2}))\to H^{2}(X_{2},\Omega_{X_{2}})=0.$
Since the isomorphism $\gamma$ is a natural one, we conclude that in that exact sequence the map
$H^{1}(X_{2},\Omega_{X_{2}})\to H^{1}(V_{2},\Omega_{X_{2}}|_{V_{2}})$
is injective (it is the inclusion $\rho$, twisted by the isomorphism $\gamma$), so that
$H^{1}(X_{2},\Omega_{X_{2}}(-V_{2}))=0.$
If we also have $H^{2}(X_{2},\Omega_{X_{2}}(-V_{2}))=0$, then we obtain the required claim. Note that since $4H_{\mathbb{P}}-2{\mathbb{E}}_{1}-2{\mathbb{E}}_{2}$ is not ample, although it is nef and big, the Akizuki-Nakano theorem [11] does not apply, see also [12].
Now consider a general quadric $S\subset{\mathbb{P}}^{4}$, such that $S\ni o_{1}$ and the strict transform $S_{1}\subset X_{1}$ contains the point $o_{2}$. Denote the strict transform of $S_{1}$ on $X_{2}$ by the symbol $S_{2}$. Obviously, the hypersurfaces $S,S_{1},S_{2}$ are non-singular and the class of divisor $S_{2}$ on $X_{2}$ is
$2H_{\mathbb{P}}-{\mathbb{E}}_{1}-{\mathbb{E}}_{2},$
so that $V_{2}\sim 2S_{2}$. Since the original quadric $S$ is non-singular, we have
$\mathop{\rm Pic}S_{2}={\mathbb{Z}}H_{S}\oplus{\mathbb{Z}}P_{1}\oplus{\mathbb{Z}}P_{2},$
where $H_{S}$ is the class of a hyperplane section and $P_{1}={\mathbb{E}}_{1}|_{S_{1}}$, $P_{2}={\mathbb{E}}_{2}|_{S_{2}}$ (recall that the pull back of a divisor is denoted by the same symbol). Repeating the arguments, performed above for the pair $X_{2}$, $V_{2}$, for the pair $X_{2}$, $S_{2}$, we obtain the exact sequence
$0\to H^{1}(X_{2},\Omega_{X_{2}}(-S_{2}))\to H^{1}(X_{2},\Omega_{X_{2}})\to H^{1}(S_{2},\Omega_{X_{2}}|_{S_{2}})\to H^{2}(X_{2},\Omega_{X_{2}}(-S_{2}))\to 0.$
However, the central map in that sequence is induced by the restriction map $\mathop{\rm Pic}X_{2}\to\mathop{\rm Pic}S_{2}$, which is an isomorphism since the quadric $S$ is smooth, so we conclude that
$H^{i}(X_{2},\Omega_{X_{2}}(-S_{2}))=0$
for $i=1,2$. Taking these equalities into account, from the short exact sequence
$0\to{\cal O}_{X_{2}}(-S_{2})\to{\cal O}_{X_{2}}\to{\cal O}_{S_{2}}\to 0,$
tensored by $\Omega_{X_{2}}(-S_{2})$, we get the isomorphism
$H^{1}(S_{2},\Omega_{X_{2}}(-S_{2})|_{S_{2}})\cong H^{2}(X_{2},\Omega_{X_{2}}(-V_{2})).$
Lemma 1.1. The space $H^{1}(S_{2},\Omega_{X_{2}}(-S_{2})|_{S_{2}})$ is trivial.
Proof is given in Subsection 1.3.
Lemma 1.1 obviously implies the claim of Theorem 0.1 for quartics of the first type.
1.3. Proof of Lemma 1.1. Our arguments are similar to the arguments of the previous subsection. Tensoring the exact sequence
$0\to{\cal O}_{S_{2}}(-S_{2})\to\Omega_{X_{2}}|_{S_{2}}\to\Omega_{S_{2}}\to 0$
by ${\cal O}_{S_{2}}(-S_{2})$ and taking into account that the class
$S_{2}|_{S_{2}}=2H_{S}-P_{1}-P_{2}$
is nef and big, we obtain the equalities
$H^{1}(S_{2},{\cal O}_{S_{2}}(-2S_{2}))=H^{2}(S_{2},{\cal O}_{S_{2}}(-2S_{2}))=0,$
so that
$H^{1}(S_{2},\Omega_{X_{2}}(-S_{2})|_{S_{2}})=H^{1}(S_{2},\Omega_{S_{2}}(-S_{2}))$
and we have to show that the space $H^{1}(S_{2},\Omega_{S_{2}}(-S_{2}))$ is trivial. Consider the free linear system $|2H_{S}-P_{1}-P_{2}|$, let $W$ be a general surface in that system. Obviously, its image in ${\mathbb{P}}^{4}$ is a non-singular del Pezzo surface of degree 4, and $W$ is its blow-up first at the point $o_{1}$, and then (of its strict transform) at the point $o_{2}$. The rank of the group $\mathop{\rm Pic}W$ is 8, and the restriction map
$\rho_{W}\colon\mathop{\rm Pic}S_{2}\to\mathop{\rm Pic}W$
is injective. For the surface $W$ the equalities $H^{i}(W,{\cal O}_{W})=0$ hold for $i=1,2$, so that $\mathop{\rm Pic}W\otimes{\mathbb{C}}=H^{1}(W,\Omega_{W})$. The injective map
$H^{1}(S_{2},\Omega_{S_{2}})\to H^{1}(W,\Omega_{W}),$
induced by the restriction of the forms, we denote by the same symbol $\rho_{W}$. Since the divisor $(2H_{S}-P_{1}-P_{2})|_{W}$ is nef and big, we get the equality $H^{1}(W,{\cal O}_{W}(-S_{2}))=0$, so that the exact sequence
$0\to{\cal O}_{W}(-S_{2})\to\Omega_{S_{2}}|_{W}\to\Omega_{W}\to 0$
implies injectivity of the map
$\beta\colon H^{1}(W,\Omega_{S_{2}}|_{W})\to H^{1}(W,\Omega_{W})$
(and also the equality $H^{0}(W,\Omega_{S_{2}}|_{W})=0$). Now let us tensor the exact sequence
$0\to{\cal O}_{S_{2}}(-S_{2})\to{\cal O}_{S_{2}}\to{\cal O}_{W}\to 0$
by $\Omega_{S_{2}}$ and obtain an exact sequence of linear maps
$0\to H^{1}(S_{2},\Omega_{S_{2}}(-S_{2}))\to H^{1}(S_{2},\Omega_{S_{2}})\stackrel{{\scriptstyle\alpha}}{{\to}}H^{1}(W,\Omega_{S_{2}}|_{W}).$
Since the composition $\beta\circ\alpha=\rho_{W}$ is injective, we conclude from here that the space $H^{1}(S_{2},\Omega_{S_{2}}(-S_{2}))=0$, as we claimed. Q.E.D. for Lemma 1.1.
1.4. Quartics of the second type. Now let us make sure that the arguments above work for quartics of the second type. Since the notations for both cases are almost identical, no changes in the formulas are needed. As we will show below (see Subsection 4.3), in the case of a second quadratic singularity $o_{2}\neq o_{1}$ the form $q_{4}$ can be arbitrary (there are some relations between $q_{3}$ and $q_{4}$, which can be satisfied at the expense of the cubic component), so that the argument in [2, Remark 1.2] works for quartics of the second type, too.
The class $4H_{\mathbb{P}}-2{\mathbb{E}}_{1}-2{\mathbb{E}}_{2}$ (and its restriction onto $V_{2}$) is still nef and big and the map $\rho$ of Subsection 1.1 is injective. The proceedings of Subsection 1.2 work without changes, except for the only one which is obvious: the quadric $S\subset{\mathbb{P}}^{4}$ simply contains the two points $o_{1}$, $o_{2}$ (and the linear system $|2H_{\mathbb{P}}-{\mathbb{E}}_{1}-{\mathbb{E}}_{2}|$ is free). Word for word the same arguments as in Subsection 1.2 reduce the factoriality of the quartic $V$ to proving Lemma 1.1, the statement of which does not change. In the arguments of Subsection 1.3 one has to change only the description of the surface $W$: now it is the strict transform of a del Pezzo surface containing the points $o_{1}$, $o_{2}$. Repeating the arguments of Subsection 1.3 word for word, we complete the proof of Theorem 0.1.
1.5. A generalization. The contents and notations of the present subsection are independent of the rest of the paper and are not used in the sequel. Let
$\pi_{i}\colon X_{i}\to X_{i-1},$
$i=1,\dots,k$, where $k\geqslant 3$, be a sequence of blow-ups of the points $o_{i-1}\in X_{i-1}$, where $X_{0}={\mathbb{P}}^{4}$, and ${\mathbb{E}}_{i}=\pi_{i}^{-1}(o_{i-1})\cong{\mathbb{P}}^{3}$ be the exceptional divisors. Let $V\subset{\mathbb{P}}^{4}$ be an irreducible hypersurface of degree 4, $V_{i}\subset X_{i}$ its strict transform, where
$V_{k}\sim 4H_{\mathbb{P}}-2\sum_{i=1}^{k}{\mathbb{E}}_{i}$
is a non-singular hypersurface. Set $Q_{i}=V_{i}\cap{\mathbb{E}}_{i}$ and assume that $Q_{i}$ is an irreducible quadric in ${\mathbb{E}}_{i}$ for all $i=1,\dots,k$.
Theorem 1.1. Assume that the divisorial class
$\Delta=2H_{\mathbb{P}}-\sum_{i=1}^{k}{\mathbb{E}}_{i}$
is nef and big and the linear system $|\Delta|$ is free (has no base points). Then the variety $V$ is factorial and
$\mathop{\rm Pic}V_{k}={\mathbb{Z}}H\bigoplus\mathop{\oplus}\limits_{i=1}^{k}{\mathbb{Z}}Q_{i}.$
Proof: repeat the arguments of Subsections 1.1 – 1.4 word for word.
2 Birational rigidity
In this section we reduce the claims of Theorems 0.2 and 0.3 to certain facts about multiplicities of a mobile linear system $\Sigma$ on the quartic $V$, which will be shown in §§3,5 for quartics of the first type and in §4 for quartics of the second type.
2.1. Maximal singularities. Up to, and including, Subsection 2.3, we consider quartics of the first type. Let $\Sigma$ be a mobile linear system on $V$ (that is, $\mathop{\rm codim}\mathop{\rm Bs}\Sigma\geqslant 2$). There is a positive integer $n\geqslant 1$ and non-negative integers $m_{1},m_{2}\in{\mathbb{Z}}_{+}$, such that
$\Sigma\subset|nH-m_{1}Q_{1}-m_{2}Q_{2}|.$
(It is easy to see that $m_{1}\geqslant m_{2}$ and $n=n[\Sigma]$ in the sense of Subsection 0.3.) Let $L_{1},\dots,L_{24}$ be the lines on $V$, passing through the point $o_{1}$. Repeating the arguments of [1, §1] (which, in their turn, repeat the arguments of [4, §§1,2]) word for word, we see that both Theorems 0.2 and 0.3 follow from a more technical claim.
Theorem 2.1. Assume that for a general divisor $D\in\Sigma$ the pair $(V,\frac{1}{n}D)$ is not canonical. Then among the 25 integers
$m_{1},\quad\mathop{\rm mult}\nolimits_{L_{1}}\Sigma,\quad\dots,\quad\mathop{\rm mult}\nolimits_{L_{24}}\Sigma$
precisely one is strictly higher than $n$, and the others are strictly smaller than $n$.
(See [1, Theorem 1.1].)
Here the word for word repeating comes to an end: the fact that the vertex $o_{1}$ of the cone $Q_{1}$ is a singular point of $V_{1}$ changes some arguments in the proof of Theorem 2.1 compared to the proof of [1, Theorem 1.1]. Let us describe in detail, precisely which cases require additional work, and which are completely similar to [1].
Let $E$ be an exceptional divisor over $V$ which is a maximal singularity of the system $\Sigma$, that is, satisfies the Noether-Fano inequality
$\mathop{\rm ord}\nolimits_{E}\Sigma>n\cdot a(E).$
The arguments of [1, Subsection 2.1] (that go back to the classical paper [5]) exclude the option that the centre of the singularity $E$ on $V$ is a non-singular point $p\neq o_{1}$. Therefore, for the centre of the singularity $E$ we have two options:
(1) this centre is the point $o_{1}$,
(2) this centre is an irreducible curve $B\subset V$.
In its turn, each of these options generates a few cases: for (1) they are
(1.1) the exceptional cone $Q_{1}$ is a maximal singularity, that is, $m_{1}>n$,
(1.2) the inequality $m_{1}\leqslant n$ holds, and the centre of $E$ on $V_{1}$ is the point $o_{2}\in Q_{1}$,
(1.3) $m_{1}\leqslant n$, and the centre of $E$ on $V_{1}$ is a point $p\in Q_{1}$, $p\neq o_{2}$,
(1.4) $m_{1}\leqslant n$, and the centre of $E$ on $V_{1}$ is a generator $R\ni o_{2}$ of the cone $Q_{1}$,
(1.5) $m_{1}\leqslant n$, and the centre of $E$ on $V_{1}$ is an irreducible curve of degree $\geqslant 2$ on $Q_{1}$;
for (2) they are
(2.1) the centre of $E$ is one of the lines $L_{1},\dots,L_{24}$,
(2.2) the centre of $E$ is a curve $B\subset V$, not containing the point $o_{1}$,
(2.3) the centre of $E$ is a curve $B$ of degree $\geqslant 2$, containing the point $o_{1}$.
In order to prove Theorem 2.1 we have to investigate all these 8 cases. Note that essentially this theorem claims that either (1.1) or (2.1) takes place, and in the latter case only one line is the centre of a maximal singularity.
2.2. Maximal curves. Assume that the centre of the maximal singularity $E$ is an irreducible curve $B\subset V$. Arguing word for word as in [1, Subsection 2.2], we exclude the case (2.2) (because in that case we may just refer to [5]) and conclude that $\mathop{\rm deg}B\leqslant 3$. If the curve $B$ is not contained in a (two-dimensional) plane, then $\mathop{\rm deg}B=3$ and $B$ is a rational normal curve in some hyperplane in ${\mathbb{P}}^{4}$.
Proposition 2.1. The following inequality holds:
$\mu=\mathop{\rm mult}\nolimits_{B}\Sigma<n.$
Proof. We argue as in [1, Subsection 2.2], but with some changes coming from the singular point $o_{2}\in V_{1}$. Let $B^{1}\subset V_{1}$ be the strict transform of $B$. If $o_{2}\not\in B^{1}$, then for a general quadric $W$ in ${\mathbb{P}}^{4}$, containing the curve $B$, we have $W^{1}\not\ni o_{2}$, so that the strict transform of the surface $S=W\cap V$, that is, $S^{1}\subset V_{1}$, is non-singular: we get the claim, similar to [1, Lemma 2.1] (where the strict transform of the surface $S$ is denoted by the symbol $S^{+}$), so that the argument, that follows immediately after the proof of [1, Lemma 2.1], can be repeated word for word, which gives the claim of our proposition.
Therefore we assume that $o_{2}\in B^{1}$. In that case for a general quadric $W\supset B$ we have $\mathop{\rm Sing}S^{1}=\{o_{2}\}$.
Lemma 2.1. The point $o_{2}$ is a non-degenerate quadratic singularity of the surface $S^{1}$, so that its strict transform $S^{2}\subset V_{2}$ is a non-singular surface, and the exceptional curve $S^{2}\cap{\mathbb{E}}_{2}$ is a non-singular conic (a plane section of the quadric $Q_{2}$).
Proof of the lemma. The strict transforms on $X_{1}$ of the quadrics in ${\mathbb{P}}^{4}$, cutting out the curve $B$, cut out $B^{1}$ on $X_{1}$. Therefore, $W_{2}\cap{\mathbb{E}}_{2}$ is a general plane in ${\mathbb{E}}_{2}\cong{\mathbb{P}}^{3}$, containing the point $B_{2}\cap{\mathbb{E}}_{2}$ (when we add the upper index 2, we mean taking the strict transform on $X_{2}$), so that
$W^{2}\cap Q_{2}=S^{2}\cap Q_{2}=(W^{2}\cap{\mathbb{E}}_{2})\cap Q_{2}$
is a section of the non-singular quadric $Q_{2}$ by a plane of general position in ${\mathbb{E}}_{2}$, containing the point $B^{2}\cap Q_{2}$. Therefore, $S^{2}\cap Q_{2}$ is a non-singular conic. Q.E.D. for the lemma.
It is easy to check that the self-intersection of the mobile part of the linear system $\Sigma^{2}|_{S^{2}}$ is
$8n^{2}-6n\mu-4\mu^{2}-(2m_{1}-m_{2})^{2}-(\mu-m_{2})^{2}\geqslant 0,$
which implies that $\mu<n$. Q.E.D. for the proposition.
(In fact, the computations performed above in the singular case and the compuations in [1, Subsection 2.2] in the non-singular case give a stronger estimate $\mu<\frac{1}{4}(-3+\sqrt{41})$, but we do not need it.)
Thus the curve $B$ is contained in some two-dimensional plane $P\ni o_{1}$ in ${\mathbb{P}}^{4}$.
The plane curve $V\cap P$ is a union of irreducible components $C_{1},\dots,C_{k}$ of degrees $d_{1},\dots,d_{k}$, where $k\leqslant 4$.
Proposition 2.2 (the plane section lemma). (i) If $d_{i}\geqslant 2$, then $\mathop{\rm mult}\nolimits_{C_{i}}\Sigma\leqslant n$.
(ii) If $d_{i}=d_{j}=1$, where $i\neq j$, then
$\mathop{\rm mult}\nolimits_{C_{i}}\Sigma+\mathop{\rm mult}\nolimits_{C_{j}}\Sigma\leqslant 2n.$
Proof will be given in §3.
Due to the presence of the singular point $o_{2}\in V_{1}$ we can not (as it was done in [1, Subsection 2.3]) simply give a reference to [4, §6]: there are cases that are not covered by the arguments in that paper, and they should be studied one by one. We put this work off until §3.
Proposition 2.2, (i) excludes the option (2.3), and the part (ii) provides the uniqueness of a maximal line in Theorem 2.1. To conclude, let us complete the proof of the claim about uniqueness in that theorem: let $L=L_{i}$, $1\leqslant i\leqslant 24$, be one of the lines and $\mu=\mathop{\rm mult}\nolimits_{L}\Sigma$. Since $L^{1}\not\ni o_{2}$, where $L^{1}\subset V_{1}$ is the strict transform of that line, we conclude that the strict transform $S^{1}\subset V_{1}$ of the section $S$ of the quartic $V$ by a general hyperplane, containing the line $L$, does not contain the point $o_{2}$, that is, $S^{1}\cap Q_{1}=S^{1}\cap{\mathbb{E}}_{1}$ is a non-singular conic. Therefore, we can repeat the claim and the proof of [1, Lemma 2.2] word for word: if one of the integers $m_{1}$, $\mu$ is strictly higher than $n$, then the other one is strictly smaller than $n$, which gives the full claim about uniqueness in Theorem 2.1.
2.3. Maximal singularities over the point $o_{1}$. Now let us consider the options (1.2)–(1.5). By the work that was done in Subsection 2.2, we may assume that the inequality $\mathop{\rm mult}\nolimits_{C}\Sigma\leqslant n$ holds for every irreducible curve $C\subset V$, and also that $m_{1}\leqslant n$.
Proposition 2.3. The case (1.2) does not take place: the centre of $E$ on $V_{1}$ is not the point $o_{2}$.
Proof. Assume the converse. Since $m_{1}\leqslant n$, the pair $(V_{1},\frac{1}{n}D^{1})$ is not canonical, where $D^{1}\in\Sigma^{1}$ is the strict transform of a general divisor $D^{1}$. Applying the inversion of adjunction (in the same way as it was done in [13] or [6]), we conclude that already $Q_{2}$ is a non-canonical singularity of that pair, that is $m_{2}>n$. But then $m_{1}\geqslant m_{2}>n$, too. Contradiction. Q.E.D. for the proposition.
Since for every line $L\subset V$, such that $L\ni o_{1}$, we have $L^{1}\not\ni o_{2}$, the arguments in [1, §3] work without any changes, so that the case (1.3) does not take place. Furthermore, the proof of [1, Lemma 2.3] works without changes, so that the case (1.5) also does not take place. However, the proof of [1, Lemma 2.4] does not work, precisely because the vertex $o_{2}$ of the cone $Q_{1}$ is a singularity of the variety $V_{1}$.
Proposition 2.4. The case (1.4) does not take place: the centre of the maximal singularity $E$ on $V_{1}$ is not a generator $R\subset Q_{1}$.
Proof will be given in §5. It requires the technique that is used in the proof of the plane section lemma, and for that reason makes an essential use of the work in §3.
This completes the proof of Theorem 2.1 for quartics of the first type.
2.4. Quartics of the second type. Now let $V$ be a quartic of the second type. Here we point out the necessary changes (mainly simplifications), that are needed in the arguments of Subsections 2.1 – 2.3. Again
$\Sigma\subset|nY-m_{1}Q_{1}-m_{2}Q_{2}|$
is a mobile linear system, however, the inetegers $m_{1}$, $m_{2}$ are not related by the ineqyality $m_{1}\geqslant m_{2}$. Let $L_{1,1},\dots,L_{1,24}$ be the lines on $V$, passing through the point $o_{1}$, and $L_{2,1},\dots,L_{2,L(o_{2})}$ the lines on $V$, passing through the point $o_{2}$. Theorem 2.1 should be replaced by a claim about $26+L(o_{2})$ multiplicities.
Theorem 2.2. Assume that for a general divisor $D\in\Sigma$ the pair $(V,\frac{1}{n}D)$ is not canonical. In that case among $26+L(o_{2})$ multiplicities
$m_{1},\quad\mathop{\rm mult}\nolimits_{L_{1,i}}\Sigma,1\leqslant i\leqslant 24,\quad m_{2},\quad\mathop{\rm mult}\nolimits_{L_{2,i}}\Sigma,1\leqslant i\leqslant L(o_{2}),$
precisely one is strictly higher than $n$, and the others are strictly smaller than $n$.
Fix a maximal singularity $E$ of the system $\Sigma$. Assume that its centre on $V$ is an irreducible curve $B\subset V$. Again $\mathop{\rm deg}B\leqslant 3$ and if $B$ is not contained in a two-dimensional plane, then $B$ is a rational normal curve in some hyperplane in ${\mathbb{P}}^{4}$. If $B$ does not contain the points $o_{1}$, $o_{2}$, then we exclude this case, just making a reference to [5]; if $B$ contains only one of them, then we exclude this case, referring to [4, 1]. It remains to check that the claim of Proposition 2.1 is true if $B$ contains both points $o_{1}$, $o_{2}$. Arguing as in the proof of Proposition 2.1, consider the surface $S\supset B$, cut out on $V$ by a quadric, such that its strict transform $S^{2}\subset V_{2}$ is non-singular and the restriction $\varphi_{1}\circ\varphi_{2}\colon S^{2}\to S$ is the blow-up of two non-degenerate quadratic singularities $o_{1}$, $o_{2}$. The restriction $\Sigma^{2}|_{S^{2}}$ has a unique fixed component, the curve $B^{2}$ of multiplicity $\mu$. The self-intersection of the mobile part of this system is
$8n^{2}-6n\mu-3\mu^{2}-m_{1}^{2}-m_{2}^{2}-(\mu-m_{1})^{2}-(\mu-m_{2})^{2}\geqslant 0,$
which excludes $B$ as the centre of the maximal singularity $E$.
The plane section lemma (Proposition 2.2) remains true.
If the plane $P$ contains only one of the two singular points, then it is sufficient to refer to the proof of the plane section lemma in [4, §6]. If $P\ni o_{1},o_{2}$, then this proof does not work for formal reasons and the claim of the plane section lemma is shown in §4.
Now in order to prove the claim about uniqueness in Theorem 2.2, it remains to show the following three facts.
Proposition 2.4. If one of the two integers $m_{1},m_{2}$ is strictly higher than $n$, then the other one is strictly smaller than $n$.
Proposition 2.5. Let $\{i,j\}=\{1,2\}$. If one of the two integers $m_{i}$ and $\mathop{\rm mult}\nolimits_{L_{j,l}}\Sigma$ is strictly higher than $n$, then the other one is strictly smaller than $n$.
Proposition 2.6. Let $L_{1,i}$ and $L_{2,j}$ be two lines on $V$, where $L_{1,i}\cap L_{2,j}=\emptyset$. Then if one of the two integers
$\mathop{\rm mult}\nolimits_{L_{1,i}}\Sigma,\quad\mathop{\rm mult}\nolimits_{L_{2,j}}\Sigma$
is strictly higher than $n$, then the other one is strictly smaller than $n$.
Proof of Proposition 2.4. Consider the section $S$ of the hypersurface $V$ by a general hyperplane, containing the line $o_{1},o_{2}]$. Since $o_{1},o_{2}]\not\subset V$, the restriction $\Sigma|_{S}$ has no fixed components, so that $\Sigma^{2}|_{S^{2}}$ does not have them, too. The self-intersection of the latter linear system is
$4n^{2}-2m_{1}^{2}-2m_{2}^{2}\geqslant 0,$
which completes the proof of the proposition.
Proof of Proposition 2.5 is put off until §4, since the point $o_{i}$ and the line $L_{j,l}$ are contained in some plane, which requires the proof of the plane section lemma.
Proof of Proposition 2.6. Set
$\mu_{1}=\mathop{\rm mult}\nolimits_{L_{1,i}}\Sigma,\quad\mu_{2}=\mathop{\rm mult}\nolimits_{L_{2,j}}\Sigma.$
Consider the section $S$ of the hypersurface $V$ by a general hyperplane, containing the line $L_{1,i}$. The restriction $\Sigma|_{S}$ has the unique fixed component $L_{1,i}$ of multiplicity $\mu_{1}$ and a base point $S\cap L_{2,j}$ of multiplicity $\mu_{2}$. Computing the self-intersection of the mobile part of the restriction $\Sigma^{1}|_{S^{1}}$ (see [1, Lemma 2.2]), we obtain the inequality
$4n^{2}-m_{1}^{2}-\mu_{1}^{2}-2n\mu_{1}-(m_{1}-\mu_{1})^{2}\geqslant\mu_{2}^{2},$
which completes the proof of the proposition.
Now in order to prove Theorem 2.2 it remains to show that if all $26+L(o_{2})$ multiplicities do not exceed $n$, then the assumption that there exists a maximal singularity $E$ leads to a contradiction. By what has already been done, the centre of $E$ on $V$ is either the point $o_{1}$, and this option is excluded in [1, §3], or the point $o_{2}$, and that option is excluded by means of the inversion of adjunction [13, 6]. Proof of Theorem 2.2 is complete. Therefore, Theorems 0.2 and 0.3 are shown for quartics of the second type.
3 The plane section lemma. I
In this section we prove the plane section lemma (Proposition 2.2) for quartics of the first type. Apart from that, we will show a certain claim about the multiplicity of the strict transform of the system $\Sigma$ on $V_{1}$ along a generator of the cone $Q_{1}$: the technique of the proof of that claim (which will be needed in §5 in the proof of Proposition 2.4) is close to the proof of the plane section lemma.
3.1. Statement of the problem. Recall that we consider quartics of the first type, so that $o_{2}$ is the vertex of the cone $Q_{1}$. Let $P\ni o_{1}$ be some two-dimensional plane, $P_{1}\subset X_{1}$ its strict transform and $R=P^{1}\cap{\mathbb{E}}_{1}$ the exceptional line of the blow up of the point $o_{1}$ on $P$. If $R$ is not a generator of the cone $Q_{1}$, then $R\not\subset Q_{1}$. Assume first that $o_{2}\not\in R$, that is, $o_{2}\not\in P^{1}$. In that case the proof of the plane section lemma in [4, §6] works without changes so that the plane section lemma is shown (see also the comments in [1, Proposition 2.1]).
Assume that $o_{2}\in P^{1}$. Set
$\mu=\mathop{\rm mult}\nolimits_{o_{1}}(P\circ V),$
where $(P\circ V)$ is the plane quartic cut out by the hypersurface $V$ on the plane $P$, it can be reducible and contain multiple components. Its affine part is given by the equation $f(z_{*})|_{P}=0$. It is easy to see that $R\subset(P_{1}\circ V^{1})$ if and only if $\mu\geqslant 3$, and $R$ comes into the effective 1-cycle $(P^{1}\circ V^{1})$ with multiplicity $\mu_{2}$.
In order to reduce the number of cases which we have to consider, assume in addition that the quartic $V$ satisfies the following condition of general position: in the notations of Subsection 0.2 the polynomial $q_{3}(z_{*})$ does not vanish identically on all generators of the cone $Q_{1}=\{q_{2}=0\}$ (quartics, for which this condition is violated, form a subset of codimension 2 in the quasi-projective variety of quartics of the first type, and moreover, for every plane $P\ni o_{1}$ the inequality $\mu\leqslant 3$ holds, that is, the effective cycle $(P\circ V)$ is neither $a_{i}L_{i}+a_{j}L_{j}$, where $a_{i}+a_{j}=4$ and $i\neq j$, nor $4L_{i}$ (here $i,j\in\{1,\dots,24\}$). Therefore, $\mu\in\{2,3\}$.
Let us list all options for the cycle $(P\circ V)$. Each option has its number in the format (Pa.b), where $a=\mu$. We consider the planes, for which $(P\circ V)$ is not an irreducible curve of degree 4, such curves can not be the centre of a maximal singularity, see Subsection 2.2.
(2.1) $(P\circ V)=C_{1}+C_{2}$, where $C_{1},C_{2}$ are irreducible conics, tangent at the point $o_{1}$, and moreover, $o_{2}\in C^{1}_{1},C^{1}_{2}$.
(2.2) $(P\circ V)=2C$, where $C$ is an irreducible conic, such that $o_{2}\in C^{1}$.
(This completes the list of options for $\mu=2$. It may seem that there are many more options, but this is not the case because of the assumption that $o_{1},o_{2}]\not\subset V$ and the condition $\mathop{\rm mult}_{o_{2}}(P^{1}\circ V_{1})=2$, taking into account that $R$ is not a generator of the cycle $(P^{1}\circ V_{1})$.)
(3.1) $(P\circ V)=C_{1}+C_{2}$, where $C_{1}$ is an irreducible cubic with the double point $o_{1}$, $C_{2}\ni o_{1}$ is a line, and moreover, $o_{2}\in C^{1}_{1}$ and $o_{2}\not\in C^{1}_{2}$. The cycle $(P^{1}\circ V_{1})$ is of the form $C_{1}^{1}+C_{2}^{1}+R$.
(3.2) $(P\circ V)=C_{1}+C_{2}+C_{3}$, where $C_{1}$ is an irreducible conic, $C_{2}$ and $C_{3}$ are lines, and moreover, $o_{2}\in C^{1}_{1}$ and $o_{2}\not\in C^{1}_{2}$, $o_{2}\not\in C^{1}_{3}$. The cycle $(P^{1}\circ V_{1})$ is of the form $C^{1}_{1}+C^{1}_{2}+C^{1}_{3}+R$.
(3.3) $(P\circ V)=C_{1}+2C_{2}$, where $C_{1}$ is an irreducible conic and $C_{2}$ is a line, $o_{2}\in C^{1}_{1}$.
The list above is complete. It is easy to see that for a quartic of the first type of general position, of the five cases listed above only (3.1) takes place for some plane $P$ (precisely because we do not need to consider planes, intersecting $V$ by an irreducible curve of degree 4). We will prove the plane section lemma for the cases (2.1), (3.1) and (3.2), in order to show that the explicit geometric version of the technique of the proof of that lemma that was used in [4] works for all cases. The quartics, for which (2.2) or (3.3) take place, form subsets of high codimension in the set of quartics of the first type.
3.2. The case of a cubic curve and a line. Assume that the case (3.1) takes place. Let $S$ be the section of $V$ by a general hyperplane, containing the plane $P$.
Lemma 3.1. The point $o_{1}$ is the only singularity of the surface $S$ and the point $o_{2}$ is the only singularity of its strict transform $S^{1}$, and the strict transform $S^{2}$ is a non-singular surface.
Proof. Consider the pencil |HP||H-P| of sections of the hypersurface $V$ by hyperplanes in ${\mathbb{P}}^{4}$, containing the plane $P$. The intersection $(S_{1}\circ S_{2})$ of two general distinct elements of that pencil is, obviously, the 1-cycle $(P\circ V)$. Since $C_{1}\neq C_{2}$ are components of multiplicity one of the latter intersection, the possible singularities of the surfaces $S_{1}$, $S_{2}$ are only the points of intersection $C_{1}\cap C_{2}$. For any point $p\in C_{1}\cap C_{2}$, which is not the point $o_{1}$, the surface $S$ has a singularity at the point $p$ if and only if $S$ is cut out by the tangent hyperplane $T_{p}V$. Therefore, for a general surface $S\in|H-P|$ we have $\mathop{\rm Sing}S=\{o_{1}\}$.
Hyperplanes in ${\mathbb{P}}^{4}$, containing the plane $P$, correspond to planes in ${\mathbb{E}}_{1}\cong{\mathbb{P}}^{3}$, containing the line $R$. Therefore we get:
$S^{1}\cap{\mathbb{E}}_{1}=S^{1}\cap E_{1}=R_{1}\cup R_{2},$
where $R_{1}=R$ and $R_{2}$ is the other generator of the cone $Q_{1}$, varying together with $S$. This is true at the scheme-theoretic level, too,
$(S^{1}\circ{\mathbb{E}}_{1})=R_{1}+R_{2},$
so that $o_{2}=R_{1}\cap R_{2}$ is the only singularity of the surface $S^{1}$.
Finally, let $P^{2}\subset X_{2}$ be the strict transform of the plane $P$. Since
$\mathop{\rm mult}\nolimits_{o_{2}}(P^{1}\circ V_{1})=2$
(the strict transform of the cubic $C_{1}$ is non-singular and $C_{2}^{1}\not\ni o_{2}$), we get
$(P^{2}\circ V_{2})=C_{1}^{2}+C_{2}^{2}+R^{2},$
that is, this scheme-theoretic intersection does not contain the exceptional line $P^{2}\cap{\mathbb{E}}_{2}$, which is the exceptional curve of the blow-up $\pi_{2}|_{P^{2}}\colon P^{2}\to P^{1}$ of the point $o_{2}$ on $P^{1}$. Thus this line $P^{2}\cap{\mathbb{E}}_{2}$ is not contained in the quadric $Q_{2}$, so that the surfaces $S^{2}$ for $S\in|H-P|$ cut out on $Q_{2}$ the pencil of conics which is cut out on $Q_{2}$ by the pencil of planes, containing the line $P^{2}\cap{\mathbb{E}}_{2}$. Therefore $S^{2}\cap Q_{2}$ is a non-singular conic, so that the surface $S^{2}$ is non-singular. Q.E.D. for the lemma.
Let $R_{3}=S^{2}\cap Q_{2}$ be the exceptional conic of the blow-up of the quadratic singularity $o_{2}\in S^{1}$. We look at the multiplicities of the linear system of curves $\Sigma_{S}=\Sigma|_{S}$, pulled back on $S^{2}$, along the curves $C_{1}^{2}$, $C_{2}^{2}=C_{2}^{1}$, $R_{1}^{2}$, $R_{2}^{2}$, $R_{3}$. Intersection numbers give the following multiplication table for these five curves:
$\left(\begin{array}[]{ccccc}-2&1&1&0&1\\ 1&-2&1&0&0\\ 1&1&-2&0&1\\ 0&0&0&-2&1\\ 1&0&1&1&-2\end{array}\right)$ (1)
(the columns and rows correspond to the order, in which these curves are listed above). The canonical class of the surfaces $S$, $S^{1}$, $S^{2}$ is equal to zero. The pull back of the hyperplane section $H_{S}$ on $S^{2}$ (denoted by the same symbol) is
$H_{S}=C_{1}^{2}+C_{2}^{2}+2R_{1}^{2}+R_{2}^{2}+2R_{3},$
the class of the restriction $E_{1,S}={\mathbb{E}}_{1}|_{S^{1}}=Q_{1}|_{S^{1}}$, pulled back on $S^{2}$, is
$E_{1,S}=R_{1}^{2}+R_{2}^{2}+R_{3},$
and $E_{2,S}={\mathbb{E}}_{2}|_{S^{2}}=Q_{2}|_{S^{2}}=R_{3}$. The class of the mobile part of the restriction of the strict transform $\Sigma^{2}$ of the linear system $\Sigma$ onto the surface $S^{2}$ is
$\xi=nH_{S}-m_{1}E_{1,S}-m_{2}E_{2,S}-bR_{1}^{2}-b_{1}C_{1}^{2}-b_{2}C_{2}^{2}=$
$(n-b_{1})C_{1}^{2}+(n-b_{2})C_{2}^{2}+(2n-m_{1}-b)R_{1}^{2}+(n-m_{1})R_{2}^{2}+(2n-m_{1}-m_{2})R_{3}.$
Clearly, this class has a non-negative intersection with each of the five curves. Now consider the $\mathbb{Q}$}-vector space
$\Xi={\mathbb{Q}}C_{1}^{2}\oplus{\mathbb{Q}}C_{2}^{2}\oplus{\mathbb{Q}}R_{1}^{2}\oplus{\mathbb{Q}}R_{2}^{2}\oplus{\mathbb{Q}}R_{3}$
and define the new classes $C_{1}^{+}$, $C_{2}^{+}$, $R_{1}^{+}$ by the formulas
$\begin{array}[]{ccl}C_{1}^{2}&=&C_{1}^{2}+\frac{1}{3}R_{2}^{2}+\frac{2}{3}R_{3},\\ C_{2}^{+}&=&C_{2}^{2},\\ R_{1}^{+}&=&R_{1}^{2}+\frac{1}{3}R_{2}^{2}+\frac{2}{3}R_{3}.\end{array}$
(Obviously, the intersection form on $\Xi$ is given by the non-degenerate matrix (1).) Set
$\Xi^{+}={\mathbb{Q}}C_{1}^{+}\oplus{\mathbb{Q}}C_{2}^{+}\oplus{\mathbb{Q}}R_{1}^{+}.$
The subspace $\Xi^{+}$ is orthogonal to the classes $R_{2}^{2}$ and $R_{3}$. The intersection form on this space is given by the matrix
$\left(\begin{array}[]{ccc}-\frac{4}{3}&1&\frac{5}{3}\\ 1&-2&1\\ \frac{5}{3}&1&-\frac{4}{3}\end{array}\right).$ (2)
Set $\xi^{+}=\xi_{1}C_{1}^{+}+\xi_{2}C_{2}^{+}+\xi_{3}R_{1}^{+}$, where $\xi_{1}=n-b_{1}$, $\xi_{2}=n-b_{2}$, $\xi_{3}=2n-m_{1}-b$.
Lemma 3.2. The intersections $(\xi^{+}\cdot C_{1}^{+})$, $(\xi^{+}\cdot C_{2}^{+})$, $(\xi^{+}\cdot R_{1}^{+})$ are non-negative.
Proof. Since $C_{1}^{+}$, $C_{2}^{+}$, $R_{1}^{+}$ are orthogonal to $R_{2}^{2}$ and $R_{3}$, we get the equalities $(\xi^{+}\cdot C_{1}^{+})=(\xi\cdot C_{1}^{+})$, $(\xi^{+}\cdot C_{2}^{+})=(\xi\cdot C_{2}^{+})$ and $(\xi^{+}\cdot R_{1}^{+})=(\xi\cdot R_{1}^{+})$. However, the right hand sides of these equalities are non-negative, because the classes $C_{1}^{+}$, $C_{2}^{+}$, $R_{1}^{+}$ are non-negative linear combinations of the classes of the standard basis of the space $\Xi$, and $\xi$ has a non-negative intersection with these classes. Q.E.D. for the lemma.
Let $\Theta=\|\theta_{ij}\|_{1\leqslant i,j\leqslant 3}$ be the inverse matrix for (2). We have the obvious equalities
$\xi_{i}=\theta_{i1}(\xi^{+}\cdot C_{1}^{+})+\theta_{i2}(\xi^{+}\cdot C_{2}^{+})+\theta_{i3}(\xi^{+}\cdot R_{1}^{+}),$
where $i=1,2,3$. Lemma 3.2 implies that if the $i$-the row (or the $i$-th column, the matrix $\Theta$ is symmetric) consists of non-negative numbers, then $\xi_{i}\geqslant 0$. It is easy to check that
$\Theta=\frac{1}{24}\left(\begin{array}[]{ccc}5&9&13\\ 9&-3&9\\ 13&9&5\end{array}\right).$
This immediately implies the inequality $\xi_{1}\geqslant 0$, which is equivalent to the estimate
$\mathop{\rm mult}\nolimits_{C_{1}}\Sigma\leqslant n.$
This is all that Proposition 2.2 claims in the case (3.1), so that the plane section lemma is shown for that case.
Note, that $\theta_{22}=-\frac{3}{24}$, and this agrees with the fact that the line $C_{2}$ can be a maximal curve of the system $\Sigma$.
In order to prove Proposition 2.2 it was sufficient to consider the orthogonal complement to the three classes $R_{1}^{2}$, $R_{2}^{2}$, $R_{3}$. We did not do that intentionally: from the explicit form of the matrix $\Theta$ we conclude that $\xi_{3}\geqslant 0$, so that the following claim is true.
Proposition 3.1. The following inequality holds
$m_{1}+\mathop{\rm mult}\nolimits_{R}\Sigma^{1}\leqslant 2n.$
Geometrically it means that the exceptional divisor of the blow-up of the infinitely near line $R$ on $V_{1}$ is not a maximal singularity of the system $\Sigma$. This fact will be needed in §5 in the proof of Proposition 2.4.
Remark 3.1. The argument above, when we restrict the linear system $\Sigma$ onto a general surface $S\in|H-P|$ and study the intersection form on the verctor space, the basis of which is formed by the numerical classes of the components of the plane section and the exceptional curves of the resolution of singularities of the surface $S$, is the explicit geometric form of the argument in [4, §6], which was carried out in terms of numerical classes of a certain blow-up of the quartic. In [4, §6] it was done in that way in order to avoid considering a large number of possible cases. In this paper we restrict the amount of cases by the conditions of general position and so it is easier to prove the plane section lemma for each case, in which the point $o_{2}$ is present, by explicit geometric arguments, resolving singularities of the surface $S$ and analysing the intersection form for a particular set of curves. However, the logic of the arguments is the same as in [4, §6]; in particular, the replacement of the classes $C_{1}^{2}$, $C_{2}^{2}$, $R_{1}^{2}$ by the classes $C_{1}^{+}$, $C_{2}^{+}$, $R_{1}^{+}$, orthogonal to the curves $R_{2}^{2}$, $R_{3}$, corresponds to the ‘‘reductions’’ in [4, §6].
For the remaining cases (2.1) and (3.2) we only point out the necessary changes in the arguments for the case (3.1) considered above.
3.3. The case of a conic and two distinct lines. Assume that the case (3.2) takes place. The claim of Lemma 3.1 remains true (the proof can be repeated word for word). On the non-singular surface $S^{2}$ in this case we have 6 curves
$C_{1}^{2},\quad C_{2}^{2}=C_{2}^{1},\quad C_{3}^{2}=C_{3}^{1},\quad R_{1}^{2},\quad R_{2}^{2},\quad R_{3},$
the intersection indices of which form the matrix
$\left(\begin{array}[]{cccccc}-2&1&1&0&0&1\\ 1&-2&0&1&0&0\\ 1&0&-2&1&0&0\\ 0&1&1&-2&0&1\\ 0&0&0&0&-2&1\\ 1&0&0&1&1&-2\end{array}\right).$
The pull back of the hyperplane section $H_{S}$ on $S^{2}$ is
$H_{S}=C_{1}^{2}+C_{2}^{2}+C_{3}^{2}+2R_{1}^{2}+R_{2}^{2}+2R_{3},$
the classes $E_{1,S}$ and $E_{2,S}$ are given by the same formulas as in the case (3.1), so that the class of the mobile part of the restriction of the strict transform of $\Sigma^{2}$ onto the surface $S^{2}$ is
$\xi=\sum_{i=1}^{3}(n-b_{i})C_{i}^{2}+(2n-m_{1}-b)R_{1}^{2}+(n-m_{1})R_{2}^{2}+(2n-m_{1}-m_{2})R_{3},$
where $b_{i}=\mathop{\rm mult}\nolimits_{C_{i}}\Sigma$ and $b=\mathop{\rm mult}\nolimits_{R}\Sigma^{1}$. By construction, $\xi$ intersects non-negatively each of the 6 curves. Consider the vector space
$\Xi={\mathbb{Q}}C_{1}^{2}\oplus{\mathbb{Q}}C_{2}^{2}\oplus{\mathbb{Q}}C_{3}^{2}\oplus{\mathbb{Q}}R_{1}^{2}\oplus{\mathbb{Q}}R_{2}^{2}\oplus{\mathbb{Q}}R_{3}$
and define the new classes $C_{1}^{+}$, $C_{2}^{+}$, $C_{3}^{+}$, $R_{1}^{+}$, orthogonal to $R_{2}^{2}$ and $R_{3}$, by the formulas
$\begin{array}[]{ccl}C_{1}^{2}&=&C_{1}^{2}+\frac{1}{3}R_{2}^{2}+\frac{2}{3}R_{3},\\ C_{2}^{+}&=&C_{2}^{2},\\ C_{3}^{+}&=&C_{3}^{2},\\ R_{1}^{+}&=&R_{1}^{2}+\frac{1}{3}R_{2}^{2}+\frac{2}{3}R_{3},\end{array}$
so that the intersection form for the basis classes of the vector space
$\Xi^{+}={\mathbb{Q}}C_{1}^{+}\oplus{\mathbb{Q}}C_{2}^{+}\oplus{\mathbb{Q}}C_{3}^{+}\oplus{\mathbb{Q}}R_{1}^{+}$
is given by the matrix
$\left(\begin{array}[]{cccc}-\frac{4}{3}&1&1&\frac{2}{3}\\ 1&-2&0&1\\ 1&0&-2&1\\ \frac{2}{3}&1&1&-\frac{4}{3}\end{array}\right).$ (3)
Set $\xi^{+}=\sum_{i=1}^{3}\xi_{i}C_{i}^{+}+\xi_{4}R_{1}^{+}$, where for $i\in\{1,2,3\}$ we have $\xi_{i}=n-b_{i}$ and $\xi_{4}=2n-m_{1}-b$. Repeating the arguments in the proof of Lemma 3.2 word for word, we conclude that $(\xi^{+}\cdot C_{i}^{+})\geqslant 0$ for $i\in\{1,2,3\}$ and $(\xi^{+}\cdot R_{1}^{+})\geqslant 0$. The inverse matrix for (3) is
$\Theta=\|\theta_{ij}\|_{1\leqslant i,j\leqslant 4}=\frac{1}{8}\left(\begin{array}[]{cccc}1&3&3&5\\ 3&-1&3&3\\ 3&3&-1&3\\ 5&3&3&1\end{array}\right).$
As in the case (3.1), the non-negativity of the numbers in the first row of the matrix $\Theta$ gives the first claim of Proposition 2.2. The equality
$\xi_{2}+\xi_{3}=\sum_{i=1}^{3}(\theta_{2i}+\theta_{3i})(\xi^{+}\cdot C_{i}^{+})+(\theta_{24}+\theta_{34})(\xi^{+}\cdot R_{1}^{+})$
and the fact that the sum of the second and third rows in the matrix $\Theta$ consists of positive numbers imply that $\xi_{2}+\xi_{3}\geqslant 0$, which is equivalent to the second claim of Proposition 2.2. Proof of the plane section lemma in the case (3.2) is complete.
The non-negativity (in fact, positivity, but the non-negativity is sufficient) of the last row of the matrix $\Theta$ gives the claim of Proposition 3.1 for the case (3.2).
3.4. The case of two conics. Assume that the case (2.1) takes place. The surface $S^{2}$ is again non-singular, and the multiplication table for the set of 5 curves $C_{1}^{2},C_{2}^{2},R_{1}^{2},R_{2}^{2},R_{3}$ is given by the matrix
$\left(\begin{array}[]{ccccc}-2&2&0&0&1\\ 2&-2&0&0&1\\ 0&0&-2&0&1\\ 0&0&0&-2&1\\ 1&1&1&1&-2\end{array}\right).$
Since in this case $\mu=2$, the exceptional line $R$ is not contained in the intersection $P^{1}\cap V_{1}$, so that for the pull back on $S^{2}$ of the hyperplane section of the surface $S$ we have
$H_{S}=C_{1}^{2}+C_{2}^{2}+R_{1}^{2}+R_{2}^{2}+2R_{3}$
(where in contrast to the previous cases $R_{1}\neq R$: the line $R\subset{\mathbb{E}}_{1}$ intersects the cone $Q_{1}$ at the only point, the vertex $o_{2}$), besides,
$E_{1,S}=R_{1}^{2}+R_{2}^{2}+R_{3}\quad\mbox{and}\quad E_{2,S}=R_{3}.$
The class of restriction of the strict transform of $\Sigma^{2}$ onto $S^{2}$ is
$\xi=\sum_{i=1}^{2}(n-b_{i})C_{i}^{2}+(n-m_{1})R_{1}^{2}+(n-m_{1})R_{2}^{2}+(2n-m_{1}-m_{2})R_{3},$
where $b_{i}$ have the same meaning as in the previous cases. The classes
$C_{1}^{+},C_{2}^{+}\in\Xi={\mathbb{Q}}C_{1}^{2}\oplus{\mathbb{Q}}C_{2}^{2}\oplus{\mathbb{Q}}R_{1}^{2}\oplus{\mathbb{Q}}R_{2}^{2}\oplus{\mathbb{Q}}R_{3}$
are defined by the equalities
$C_{i}^{+}=C_{i}^{2}+\frac{1}{2}R_{1}^{2}+\frac{1}{2}R_{2}^{2}+R_{3},$
where $i=1,2$, and the intersection form on $\Xi^{+}={\mathbb{Q}}C_{1}^{+}\oplus{\mathbb{Q}}C_{2}^{+}$ is given by the matrix
$\left(\begin{array}[]{cc}-1&3\\ 3&-1\end{array}\right),$
the inverse of which is
$\Theta=\frac{1}{8}\left(\begin{array}[]{cc}1&3\\ 3&1\end{array}\right).$
Since all coefficients of that matrix are positive, the plane section lemma is shown for the case (2.1). Q.E.D. for Proposition 2.2.
3.5. The case of a quartic with a triple point. Assume that $P\cap V$ is an irreducible curve $C$ of degree 4, and moreover, $\mathop{\rm mult}\nolimits_{o_{1}}C=3$. Although this case is not needed for the proof of the plane section lemma (we know that the curve $C$ can not be maximal), it is needed in order to complete the proof of Proposition 3.1: obviously,
$(P^{1}\circ V_{1})=C^{1}+R,$
where $R$ is a generator of the cone $Q_{1}$, here $o_{2}\in C^{1}$ is a smooth point (the curve $C^{1}$ is non-singular), so that $C^{1}$ meets $R$ in three points, one of which is $o_{2}$. For a general surface $S\in|H-P|$ its strict transform $S^{2}$ is non-singular, the multiplication table for the set of curves $C^{2}$, $R_{1}^{2}$, $R_{2}^{2}$, $R_{3}$ (where $R=R_{1}$) is given by the matrix
$\left(\begin{array}[]{cccc}-2&2&0&1\\ 2&-2&0&1\\ 0&0&-2&1\\ 1&1&1&-2\end{array}\right).$
We have the following equalities:
$\begin{array}[]{ccl}H_{S}&=&C^{2}+2R_{1}^{2}+R_{2}^{2}+2R_{3},\\ E_{1,S}&=&R_{1}^{2}+R_{2}^{2}+R_{3},\\ E_{2,S}&=&R_{3}.\end{array}$
The classes
$C^{+},R_{1}^{+}\in\Xi={\mathbb{Q}}C^{2}\oplus{\mathbb{Q}}R_{1}^{2}\oplus{\mathbb{Q}}R_{2}^{2}\oplus{\mathbb{Q}}R_{3}$
are given by the equalities
$\begin{array}[]{ccl}C^{+}&=&C^{2}+\frac{1}{3}R_{2}^{2}+\frac{2}{3}R_{3},\\ R_{1}^{+}&=&R_{1}^{2}+\frac{1}{3}R_{2}^{2}+\frac{2}{3}R_{3},\end{array}$
the intersection form on $\Xi^{+}={\mathbb{Q}}C^{+}\oplus{\mathbb{Q}}R_{1}^{+}$ is given by the matrix
$\frac{1}{3}\left(\begin{array}[]{cc}-4&8\\ 8&-4\end{array}\right),$
the inverse for which is
$\Theta=\frac{1}{4}\left(\begin{array}[]{cc}1&2\\ 2&1\end{array}\right).$
The non-negativity of coefficients of that matrix implies Proposition 3.1 for the case under consideration.
4 The plane section lemma. II
In this section we prove the plane section lemma (Proposition 2.2) for quartics of the second type. After that we prove Proposition 0.1: for a quartics of the second type of general position there are 24 distinct lines through the point $o_{2}$, lying on the quartic.
4.1. Statement of the problem. We consider quartics of the second type with two singular points: $o_{1}$ of rank 3 (which is resolved by one blow-up) and $o_{2}$ of rank 4. If the plane $P$ contains only one of these two points, then the proof of the plane section lemma given in [4, §6], works word for word. For this reason, assume that $P\ni o_{1},o_{2}$.
We know that the line $o_{1},o_{2}]\not\subset V$. Assume in addition that for every plane $P\supset[o_{1},o_{2}]$ the effective 1-cycle $(P\circ V)$ does not contain multiple components (quartics of the second type, for which this condition is not satisfied, form a subset of positive codimension in the set of quartics of the second type). Since $(P\circ V)=P\cap V$ is singular at the points $o_{1}$, $o_{2}$, does not contain the line $o_{1},o_{2}]$ and is reducible (as we know, an irreducible curve of degree $\geqslant 4$ can not be maximal for $\Sigma$), the following cases are possible:
(1) $P\cap V$ is the union of four distinct lines $C_{1}$, $C_{2}$, $C_{3}$, $C_{4}$, where $o_{1}\in C_{1},C_{2}$ and $o_{2}\in C_{3},C_{4}$;
(2.1) $(P\circ V)=C_{1}+C_{2}+C_{3}$, where $C_{1}$ is an irreducible conic, $C_{1}\ni o_{1},o_{2}$, $C_{2}$ and $C_{3}$ are lines, containing, respectively, $o_{1}$ and $o_{2}$;
(2.2) $(P\circ V)=C_{1}+C_{2}$, where $C_{1}$, $C_{2}$ are irreducible conics, passing through both points $o_{1}$, $o_{2}$;
(3) $(P\circ V)=C_{1}+C_{2}$, where $C_{1}$ is an irreducible cubic with a double point and $C_{2}$ is aline; in principle we should consider separately the cases (3.1), when $o_{1}\in C_{1}$ is a double point and $o_{2}\in C_{1}\cap C_{2}$, and (3.2), when $o_{2}\in C_{1}$ is a double point and $o_{1}\in C_{1}\cap C_{2}$, however, the arguments are identical in both cases and we will consider the case (3.1) only.
In the next subsection we prove Proposition 2.2, using the technique that was applied in §3. We will just point out the necessary changes in the arguments and computations of the previous section.
4.2. Proof of the plane section lemma. In the notations of Subsection 3.2, let $S\in|H-P|$ be a general surface. Since the cycle $(P\circ V)$ has no multiple components, we can argue as in the proof of Lemma 3.1 and conclude that
$\mathop{\rm Sing}S=\{o_{1},o_{2}\},$
and the strict transform $S^{2}\subset V_{2}$ is non-singular. Obviously, $S^{2}\cap{\mathbb{E}}_{1}$ and $S^{2}\cap{\mathbb{E}}_{2}$ are non-singular conics, we denote them by the symbols $R_{1}$ and $R_{2}$.
The case of four lines. Assume that the case (1) takes place. The multiplication table for the six curves
$R_{1},\quad C_{1}^{2},\quad C_{2}^{2},\quad C_{3}^{2},\quad C_{4}^{2},\quad R_{2}$
(this order is more convenient) is given by the ($\times6$)-matrix
$\left(\begin{array}[]{cccccc}-2&1&1&0&0&0\\ 1&-2&0&1&1&0\\ 1&0&-2&1&1&0\\ 0&1&1&-2&0&1\\ 0&1&1&0&-2&1\\ 0&0&0&1&1&-2\end{array}\right).$
Introducing the classes
$C_{1}^{+}=C_{1}^{2}+\frac{1}{2}R_{1},\quad C_{2}^{+}=C_{2}^{2}+\frac{1}{2}R_{1},$
$C_{3}^{+}=C_{3}^{2}+\frac{1}{2}R_{2},\quad C_{4}^{+}=C_{4}^{2}+\frac{1}{2}R_{2},$
we obtain for $C_{1}^{+}$, $C_{2}^{+}$, $C_{3}^{+}$, $C_{4}^{+}$ the multiplication table
$\frac{1}{2}\left(\begin{array}[]{cccc}-3&1&2&2\\ 1&-3&2&2\\ 2&2&-3&1\\ 2&2&1&-3\end{array}\right).$
The inverse matrix is
$\Theta=\frac{1}{12}\left(\begin{array}[]{cccc}-1&5&4&4\\ 5&-1&4&4\\ 4&4&-1&5\\ 4&4&5&-1\end{array}\right).$
Since the sum of any two distinct rows of the matrix $\Theta$ is a vector, all coordinates of which are positive, we obtain the claim (ii) of Proposition 2.2 (the claim (i) in this case does not apply).
The case of a conic and two lines. Assume that the case (2.1) takes place. Here the multiplication table for $C_{1}^{+}$, $C_{2}^{+}$, $C_{3}^{+}$ is given by the matrix
$\frac{1}{2}\left(\begin{array}[]{ccc}-2&3&3\\ 3&-3&2\\ 3&2&-3\end{array}\right),$
the inverse of which is
$\Theta=\frac{1}{40}\left(\begin{array}[]{ccc}5&15&15\\ 15&-3&13\\ 15&13&-3\end{array}\right),$
which implies both claims of Proposition 2.2 in the case under consideration.
The case of two conics. Assume that the case (2.2) takes place. Here the multiplication table for
$C_{1}^{+}=C_{1}^{2}+\frac{1}{2}R_{1}+\frac{1}{2}R_{2},\quad C_{2}^{+}=C_{2}^{2}+\frac{1}{2}R_{1}+\frac{1}{2}R_{2}$
is given by the matrix
$\left(\begin{array}[]{cc}-1&3\\ 3&-1\end{array}\right),$
the inverse of which is
$\Theta=\frac{1}{8}\left(\begin{array}[]{cc}1&3\\ 3&1\end{array}\right),$
which proves the plane section lemma in this case.
The case of a cubic curve and a line. Assume that the case (3.1) takes place. Here the multiplication table for $C_{1}^{+}=C_{1}^{2}+R_{1}+\frac{1}{2}R_{2}$, $C_{2}^{+}=C_{2}^{2}+\frac{1}{2}R_{2}$, is given by the matrix
$\frac{1}{2}\left(\begin{array}[]{cc}1&5\\ 5&-3\end{array}\right),$
the inverse of which is
$\Theta=\frac{1}{14}\left(\begin{array}[]{cc}3&5\\ 5&-1\end{array}\right).$
This completes the proof of the plane section lemma for quartics of the second type.
4.3. Proof of Proposition 2.5. Let us show that one of the singular points $o_{i}$, $i\in\{1,2\}$, and a line $L$, passing through the other singular point, can not maximal simultaneously. We will show the inequality
$m_{i}+\mathop{\rm mult}\nolimits_{L}\Sigma\leqslant 2n,$
which immediately implies Proposition 2.5. Let $P=\langle o_{i},L\rangle$ be the plane, containing this point and this line. Proof follows the scheme of the proof of the plane section lemma in the cases (1), (2.1) and (3), see Subsection 4.1, we only need a different orthogonal projection. Let us assume that $i=1$.
The case of four lines. The class of the free part of the restriction of the system $\Sigma^{2}$ onto the surface $S^{2}$ in this case is
$\xi=nH_{S}-m_{1}E_{S,1}-m_{2}E_{S,2}-\sum_{j=1}^{4}\mu_{j}C_{j}^{2}.$
Since $H_{S}=R_{1}+R_{2}+\sum_{j=1}^{4}C_{j}^{2}$, $E_{S,1}=R_{1}$, $E_{S,2}=R_{2}$, this class can be re-written in the form
$\xi=(n-m_{1})R_{1}+(n-m_{2})R_{2}+\sum_{j=1}^{4}(n-\mu_{j})C_{j}^{2}.$
Since we are interested in the multiplicities $m_{1}$ and $\mu_{j}$, $j\in\{3,4\}$ only, let us consider in
$\Xi={\mathbb{Q}}R_{1}\oplus{\mathbb{Q}}R_{2}\bigoplus\mathop{\oplus}\limits_{j=1}^{4}{\mathbb{Q}}C_{j}^{2}$
the orthogonal complement to $C_{1}^{2}$, $C_{2}^{2}$, $R_{2}$:
$\begin{array}[]{ccl}R_{1}^{+}&=&R_{1}+\frac{1}{2}C_{1}^{2}+\frac{1}{2}C_{2}^{2},\\ C_{3}^{+}&=&C_{3}^{2}+\frac{1}{2}C_{1}^{2}+\frac{1}{2}C_{2}^{2}+\frac{1}{2}R_{2},\\ C_{4}^{+}&=&C_{4}^{2}+\frac{1}{2}C_{1}^{2}+\frac{1}{2}C_{2}^{2}+\frac{1}{2}R_{2}.\end{array}$
The multiplication table for these three classes is given by the matrix
$\frac{1}{2}\left(\begin{array}[]{ccc}-2&2&2\\ 2&-1&3\\ 2&3&-1\end{array}\right),$
the inverse matrix for which is
$\frac{1}{12}\left(\begin{array}[]{ccc}-4&4&4\\ 4&-1&5\\ 4&5&-1\end{array}\right).$
Since the sum of the first and the second (or the second and the third) rows is non-negative, this means precisely that
$(n-m_{1})+(n-\mu_{3})\geqslant 0$
(respectively, $(n-m_{1})+(n-\mu_{4})\geqslant 0$). Proof of Proposition 2.5 for the case of four lines is complete.
The case of a conic and two lines. Here $L=C_{3}$, and the multiplication table for the five curves $R_{1}^{2},C_{1}^{2},C_{2}^{2},C_{3}^{2},R_{2}^{2}$ is given by the matrix (we did not give it in Subsection 4.2):
$\left(\begin{array}[]{ccccc}-2&1&1&0&0\\ 1&-2&1&1&1\\ 1&1&-2&1&0\\ 0&1&1&-2&1\\ 0&1&0&1&-1\end{array}\right).$
The orthogonal complement to $C_{1}^{2}$, $C_{2}^{2}$ is given by the formulas
$\begin{array}[]{ccl}R_{1}^{+}&=&R_{1}^{2}+C_{1}^{2}+C_{2}^{2},\\ C_{3}^{+}&=&C_{3}^{2}+C_{1}^{2}+C_{2}^{2},\\ R_{2}^{+}&=&R_{2}^{2}+\frac{2}{3}C_{1}^{2}+\frac{1}{2}C_{2}^{2}.\end{array}$
The multiplication table for these classes is
$\left(\begin{array}[]{ccc}0&2&1\\ 2&0&2\\ 1&2&-\frac{4}{3}\end{array}\right),$
and the inverse matrix is
$\frac{1}{40}\left(\begin{array}[]{ccc}-12&14&12\\ 14&-3&6\\ 12&6&-12\end{array}\right).$
Since the sum of the first and second rows consists of non-negative numbers, Proposition 2.5 is shown in this case.
The case of a cubic curve and a line. Here the multiplication table for $R_{1}^{2},C_{1}^{2},C_{2}^{2},R_{2}^{2}$ (where $C_{1}$ is the plane cubic and $C_{2}$ is the line that we need) is
$\left(\begin{array}[]{cccc}-2&2&0&0\\ 2&-2&2&1\\ 0&2&-2&1\\ 0&1&1&-2\end{array}\right).$
The orthogonal complement to $C_{1}^{2}$, $R_{2}^{2}$ is given by the formulas
$\begin{array}[]{ccl}R_{1}^{+}&=&R_{1}^{2}+\frac{4}{3}C_{1}^{2}+\frac{2}{3}R_{2}^{2},\\ C_{2}^{+}&=&C_{2}^{2}+\frac{5}{3}C_{1}^{2}+\frac{4}{3}R_{2}^{2},\end{array}$
the multiplication table for $R_{1}^{+}$, $C_{2}^{+}$ is
$\frac{1}{3}\left(\begin{array}[]{cc}2&10\\ 10&8\end{array}\right),$
and the inverse matrix is
$\frac{1}{14}\left(\begin{array}[]{cc}-4&5\\ 5&-1\end{array}\right).$
Q.E.D. for Proposition 2.5.
4.4. Lines through the point $o_{2}$. Let us show Proposition 0.1. In the coordinate notations of Subsection 0.2, using a suitable linear change of coordinates, we assume that $o_{2}=(0,0,1,0)$. The form $q_{2}(z_{1},z_{2},z_{3})=z_{1}^{2}+z_{2}^{2}+z_{3}^{2}$ is fixed, $q_{2}(o_{2})\neq 0$, since the line $o_{1},o_{2}]$ can not be contained in the tangent cone $T_{o_{1}}V$ (otherwise, $o_{1},o_{2}]\subset V$, contrary to our assumption). Let us find out, what conditions for the polynomials $q_{3}(z_{*})$ and $q_{4}(z_{*})$ are imposed by the requirement $o_{2}\in\mathop{\rm Sing}V$. We will need some routine coordinate computations, which we will reduce to the minimum, only presenting the results of each step. Write down
$q_{3}(z_{*})=\sum_{i=0}^{3}z_{3}^{3-i}q_{3,i},\quad q_{4}(z_{*})=\sum_{i=0}^{4}z_{3}^{4-i}q_{4,i},$
where $q_{3,i}$ and $q_{4,i}$ are homogeneous polynomials of degree $i$ in $z_{1}$, $z_{2}$, $z_{4}$. Since the blow-up of the point $o_{1}$ resolves that singularity, $q_{3,3}(z_{1},z_{2},z_{4})$ contains $z_{4}^{3}$ with a non-zero coefficient. Consider the new coordinate system $(z_{1},z_{2},u,z_{4})$, where $z_{3}=1+u$.
Lemma 4.1. The condition $o_{2}\in V$ is equivalent to the equality $q_{3,0}+q_{4,0}=-1$.
Proof: obvious computations.
Assuming the equality of the lemma, let us consider the condition $o_{2}\in\mathop{\rm Sing}V$.
Lemma 4.2. The point $o_{2}$ is a singularity of the hypersurface $V$ if and only if the the following equalities hold:
$3q_{3,0}+4q_{4,0}=-2,\quad q_{3,1}+q_{4,1}\equiv 0.$
Proof: obvious computations.
From these lemmas we get the equalities $q_{3,0}=-2$, $q_{4,0}=1$. We also assume that the second equality of Lemma 4.2 holds.
Lemma 4.3 In the coordinates $(z_{1},z_{2},u,z_{4})$ the quadratic part of the polynomial $f$ at the point $o_{2}$ is
$w_{2}=u^{2}+uq_{4,1}+(z_{1}^{2}+z_{2}^{2}+q_{3,2}+q_{4,2}),$ (4)
the cubic part of the polynomial $f$ at the point $o_{2}$ is
$w_{3}=2u^{3}+2u^{2}q_{4,1}+u(q_{3,2}+2q_{4,2})+(q_{3,3}+q_{4,3}),$
and the component of degree 4 is
$q_{4}(z_{1},z_{2},u,z_{4})=\sum_{i=0}^{4}u^{4-i}q_{4,i}(z_{1},z_{2},z_{4}).$
Proof: obvious computations.
It is clear from the explicit formula (4), that for a general quartic of the second type the rank of the point $o_{2}$ is equal to 4. Identifying
${\mathbb{E}}_{2}={\mathbb{P}}_{(z_{1}:z_{2}:u:z_{4})}^{3},$
we see that the quadric $Q_{2}$ is the set of zeros of the polynomial (4), where the point $p=(0:0:1:0)$ does not lie on $Q_{2}$. Furthermore, for the fixed forms $q_{4,1}$, $q_{3,2}$, $q_{4,2}$ the polynomial $w_{3}$ contains an arbitrary cubic polynomial $(q_{3,3}+q_{4,3})$ in $(z_{1},z_{2},z_{4})$, that is, the equation of an arbitrary cubic cone with the vertex at the point $p$, which implies that for a general quartic of the second type the equation $w_{3}|_{Q_{2}}=0$ defines on the quadric a non-singular curve of bi-degree (3,3). Finally the polynomial $q_{4}(z_{1},z_{2},u,z_{4})$ contains an arbitrary component $q_{4,4}(z_{1},z_{2},z_{4})$ (that is, the equation of an arbitrary cone of degree 4 with the vertex at the point $p$), which is not contained in the equation, defining $Q_{2}$ and the curve $\{w_{3}|_{Q_{2}}=0\}$. It follows that the system of equations
$w_{3}|_{Q_{2}}=0,\quad q_{4}(z_{1},z_{2},u,z_{4})|_{Q_{2}}=0$
has precisely 24 distinct solutions. Q.E.D. for Proposition 0.1.
5 Infinitely near lines
In this section we consider only quartics of the first type. We show Proposition 2.4.
5.1. The technique of counting multiplicities. Assume (contrary to the claim of Proposition 2.4) that the centre of the maximal singularity $E$ on $V_{1}$ is a generator $R$ of the cone $Q_{1}$. By Proposition 3.1 the singularity $E$ is not the exceptional divisor of the blow-up of $R$. Let us re-denote the strict transform $R^{2}$ by the symbol $B_{1}$, it is a non-singular rational curve on the non-singular three-dimensional variety $V_{2}$. Let
$\sigma_{i}\colon V^{(i)}\to V^{(i-1)}$
be the sequence of blow-ups of curves $B_{i-1}\subset V^{(i-1)}$ with the exceptional divisors $E^{(i)}\subset V^{(i)}$, $E^{(i)}=\sigma_{i}^{(-1)}(B_{i-1})$, where $i=2,\dots,k+1$, $k\geqslant 2$, such that:
$V^{(1)}=V_{2}$ and $B_{1}=R^{2}$,
— the centre $B_{i-1}$ of each of these blow-ups is the centre of the maximal singularity $E$ on $V^{(i-1)}$,
$E^{(k+1)}$ realizes $E$ on $V^{(k+1)}$ (in other words, $E^{(k+1)}$ is the centre of $E$ on $V^{(k+1)}$).
Thus this sequence of blow-ups is the resolution of the maximal singularity $E$ in the sense of [15, Chapter 2]. The inequality $k\geqslant 2$ holds by Proposition 3.1. Let $\Sigma^{(i)}$ be the strict transform of the system $\Sigma$ on $V^{(i)}$. Set
$\nu_{i}=\mathop{\rm mult}\nolimits_{B_{i-1}}\Sigma^{(i-1)},$
$i=2,\dots,k+1$, then the Noether-Fano inequality for $E$ (that is, the condition that $E$ is a maximal singularity of the system $\Sigma$) takes the form
$p_{1}m_{1}+\sum_{i=2}^{k+1}p_{i}\nu_{i}>n\left(p_{1}+\sum_{i=2}^{k+1}p_{i}\right),$
where $p_{i}$ is the number of paths in the graph $\Gamma$ of the singularity $E=E^{(k+1)}$ from the vertex $(k+1)$ to the vertex $i$, see [15, Chapter 2] and [14]. We will not dwell on the details, because in the case under consideration this formalism simplifies essentially: one may assume that $\nu_{k+1}>n$, whence by the inequality $m_{1}\leqslant n$ it follows that the graph $\Gamma$ is a chain, all $p_{i}$ are equal to 1, for $i\leqslant k$ the curve $B_{i}\subset E^{(i)}$ is a section of the projection $E^{(i)}\to B_{i-1}$, so that all curves $B_{i}$ and all varieties $V^{(i)}$ are non-singular. The Noether-Fano inequality takes the simplified form
$m_{1}+\nu_{2}+\dots\nu_{k+1}>(k+1)n.$
Let $Z=(D_{1}\circ D_{2})$ be the self-intersection of the system $\Sigma$, where $D_{1},D_{2}\in\Sigma$ are general divisors. This is and effective 1-cycle on $V$, for which, according to the technique of counting multiplicities, see [15, Chapter 2, Proposition 2.4, Corollary 2.5], and also [15, Chapter 2, Subsection 2.2], [1, §3], we have the estimate
$\mathop{\rm mult}\nolimits_{o_{1}}Z>\frac{2(k+1)^{2}}{2k+1}n^{2}.$
It is easy to check that the right hand side of this inequality is strictly higher than $4n^{2}$ for $k\geqslant 3$. Since $\mathop{\rm deg}Z=4n^{2}$, we conclude that $k=2$. Thus the maximal singularity is realized by the blow-up of the curve $B_{2}$ (a section of the ruled surface $E^{(2)}\to B_{1}=R^{2}$), and the Noether-Fano inequality takes the form of the estimate
$m_{1}+\nu_{2}+\nu_{3}>3n.$ (5)
However, for this (only this) case $k=2$ the standard technique of counting multiplicities does not make it possible to exclude the maximal singularity, and we need to study this case in more detail.
5.2. Curves on the quadric $Q_{2}$. Let $D\in\Sigma$ be a general divisor, $D^{2}$ its strict transform on $V_{2}$. Since $D^{2}\sim nH-m_{1}Q_{1}-m_{2}Q_{2}$, the restriction $D^{2}|_{Q_{2}}$ is an effective 1-cycle on $Q_{2}\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ of bi-degree $(m_{2},m_{2})$ (in other words, this 1-cycle is cut out on $Q_{2}$ by a surface of degree $m_{2}$ in ${\mathbb{E}}_{2}\cong{\mathbb{P}}^{3}$).
Proposition 5.1. The following estimate is true:
$\nu_{2}+\nu_{3}\leqslant 3m_{2}.$ (6)
Proof. Set
$p_{1}=B_{1}\cap Q_{2}$
(the intersection is transversal, since $B_{1}=R^{2}$ intersects ${\mathbb{E}}_{2}$ transversally at this point) and let $N=\sigma_{2}^{-1}(p_{1})\subset E^{(2)}$ be the fibre of the ruled surface $E^{(2)}$ over the point $p_{1}\in B_{1}$. Obviously, $Q_{2}^{(2)}$ (the strict transform of $Q_{2}$ on $V^{(2)}$) is the result of blowing up the point $p_{1}\in Q_{2}$, and $N=Q_{2}^{(2)}\cap E^{(2)}$ is the exceptional line of that blow-up. Furthermore, the restriction of the divisor
$D^{(2)}\sim nH-m_{1}Q_{1}-m_{2}Q_{2}-\nu_{2}E^{(2)}$
onto the surface $Q_{2}^{(2)}$ is precisely
$\left(\sigma_{2}|_{Q_{2}^{(2)}}\right)^{*}D^{2}|_{Q_{2}}-\nu_{2}N.$
This effective 1-cycle on $Q_{2}^{(2)}$ has multiplicity $\geqslant\nu_{2}$ at the point $p_{2}=B_{2}\cap Q_{2}^{(2)}=B_{2}\cap N$ (the inetrsections are transversal, since $B_{2}$ is a section of the ruled surface $E^{(2)}/B_{1}$).
Therefore, we obtain the following geometric situation. On the non-singular quadric $Q_{2}\subset{\mathbb{E}}_{2}$ there is an effective 1-cycle $Y$ of bi-degree $(m_{2},m_{2})$, satisfying the inequality $\mathop{\rm mult}\nolimits_{p_{1}}Y\geqslant\nu_{2}$. On the blow-up of the point $p_{1}\in Q_{2}$ the effective 1-cycle $Y-\nu_{2}N$ (recall that $N$ is the exceptional line of that blow-up, and the pull back of $Y$, as a divisor, from $Q_{2}$ onto that blow-up is denoted by the same symbol) has multiplicity $\geqslant\nu_{3}$ at the point $p_{2}\in N$. We will consider the point $p_{2}$ as a point, infinitely near to $p_{1}$, in the sense of the blow-up of the point $p_{1}$ on ${\mathbb{E}}_{2}\cong{\mathbb{P}}^{3}$, and let $p_{1},p_{2}]\subset{\mathbb{E}}_{2}$ be the line on ${\mathbb{E}}_{2}$, passing through $p_{1}$ in the direction of $p_{2}$. There are two options:
(1) $p_{1},p_{2}]\not\subset Q_{2}$, and then $p_{1},p_{2}]\cap Q_{2}=\{p_{1}\}$ (set-theoretically),
(2) $p_{1},p_{2}]\subset Q_{2}$ is one of the two lines on $Q_{2}$, passing through the point $p_{1}$.
Assume that the case (1) takes place. Let $\Lambda\subset Q_{2}$ be the section of $Q_{2}$ by a general plane in ${\mathbb{E}}_{2}$, containing the line $p_{1},p_{2}]$. Then the intersection of $\Lambda$ with the support of the cycle $Y$ is zero-dimensional, so that we get:
$2m_{2}=(\Lambda\cdot Y)\geqslant(\Lambda\cdot Y)_{p_{1}}\geqslant\nu_{2}+\nu_{3},$
which is considerably stronger than the inequality of Proposition 5.1.
Assume that the case (2) takes place. Write down $Y=a[p_{1},p_{2}]+Y^{+}$, where $a\in{\mathbb{Z}}_{+}$ and the effective 1-cycle $Y^{+}$ of bi-degree $(m_{2}-a,m_{2})$ does not contain the line $p_{1},p_{2}]$ as a component. We get
$m_{2}=([p_{1},p_{2}]\cdot Y^{+})\geqslant\mathop{\rm mult}\nolimits_{p_{1}}Y^{+}+\mathop{\rm mult}\nolimits_{p_{2}}(Y^{+})^{(2)},$
where the symbol $(Y^{+})^{(2)}$ means the strict transform of the 1-cycle $Y^{+}$ on the blow-up of the point $p_{1}$ on $Q_{2}$ (or on $V^{(2)}$, which is the same). Now we get:
$\nu_{2}+\nu_{3}\leqslant\mathop{\rm mult}\nolimits_{p_{1}}Y+\mathop{\rm mult}\nolimits_{p_{2}}Y^{(2)}=\mathop{\rm mult}\nolimits_{p_{1}}Y^{+}+\mathop{\rm mult}\nolimits_{p_{2}}(Y^{+})^{(2)}+2a\leqslant 3m_{2},$
where $Y^{(2)}$ means the strict transform. This completes the proof of Proposition 5.1.
5.3. The two techniques of the proof. For a quartic of the first type of general position there are only two options, see Subsections 3.1 and 3.5:
(1) the case of general position, when $(V\circ P)=V\cap P$ is an irreducible curve $C$ of degree 4 with the triple point $o_{1}$, so that $o_{2}\in C^{1}$ is a non-singular point (this case takes place for allmost all generators $R$, except for finitely many of them),
(2) $(V\circ P)=C+L$, where $C$ is an irreducible cubic curve with the double point $o_{1}$, $L\ni o_{1}$ is a line, and moreover, $o_{2}\not\in L^{1}$ and $o_{2}\in C^{1}$ is a non-singular point.
We will exclude these two cases in different ways: the method of excluding the case (1) does not work in the case (2), and the method of excluding the case (2) does not apply to finitely many generators $R$ and planes $P$, for which the curve $C$ has at the point $o_{1}$ two branches, a non-singular one and a cuspidal one, and moreover, the point $o_{2}$ lies exactly on the cuspidal branch. The case (1) will be excluded by means of the ‘‘quadratic’’ technique, which is more traditional and simpler. Our approach to the proof of Proposition 2.4 in the case (1) is to estimate the input of the curve $C$ to the self-intersection $Z$ of the system $\Sigma$; we will show that this input is sufficiently big, which will make it possible to obtain a contradiction by means of the standard technique of counting multiplicities:
$\mathop{\rm mult}\nolimits_{o_{1}}Z\geqslant 2m_{1}^{2}+\nu_{2}^{2}+\nu_{3}^{2},$ (7)
see [1, §3], although it does not give the $4n^{2}$-inequality.
In the case (2) (that is, for finitely many generators $R$ of the cone $Q_{1}$) this approach does not give a contradiction, and we will argue in a different way: using the linear technique of §3, we will estimate the multiplicity of the curve $C$ with respect to the system $\Sigma$ and make the estimate of Proposition 5.1 stronger, which will lead to a contradiction without the technique of counting multiplicities.
We start with the case of general position.
5.4. The case of an irreducible plane quartic. We consider the situation of Subsection 3.5. In the notations of Subsection 3.5 we have on the non-singular surface $S^{2}$ (where $S\in|H-P|$ is a general element of the pencil):
$H_{S}=C^{2}+2R_{1}^{2}+R_{2}^{2}+2R_{3}$
is the pull back of the hyperplane section,
$E_{1,S}=R_{1}^{2}+R_{2}^{2}+R_{3},\quad E_{2,S}=R_{3}.$
The class of the mobile part of the restriction of the system $\Sigma^{2}$ onto $S^{2}$ is
$\xi=nH_{S}-m_{1}E_{1,S}-m_{2}E_{2,S}-bR_{1}^{2}-b_{C}C^{2},$
where $b_{C}=\mathop{\rm mult}\nolimits_{C}\Sigma$, and the other coefficients have the same meaning as in §3; of course, $b=\nu_{2}$. Re-writing
$\xi=(n-b_{C})C^{2}+(2n-m_{1}-b)R_{1}^{2}+(n-m_{1})R_{2}^{2}+(2n-m_{1}-m_{2})R_{3}$
and multiplying by $C^{2}$, we get (see Subsection 3.5):
$b_{C}\geqslant\frac{1}{2}(3m_{1}+m_{2}+2\nu_{2}-4n).$ (8)
(from the inequalities (5) and (6) it follows that the right hand side is strictly positive). It is easy to check that for a fixed value $\nu_{2}+\nu_{3}=2\theta$ the estimates (7) and (8) are the worst for $\nu_{2}=\nu_{3}=\theta$ (recall that $\nu_{2}\geqslant\nu_{3}$: on a non-singular variety multiplicities do not increase), so that we may assume that this equality takes place.
Consider first the case $\theta\in(n,\frac{9}{8}n]$. In that case the inequality
$3n-2\theta\geqslant\frac{2}{3}\theta$
holds, so that, making the estimates (7) and (8) worse, set $m_{1}=3n-2\theta$ and $m_{2}=\frac{2}{3}\theta$, whence we get:
$b_{C}\geqslant\frac{1}{6}(15n-10\theta)$
and
$\mathop{\rm mult}\nolimits_{o_{1}}Z\geqslant 18n^{2}-24n\theta+10\theta^{2}.$
Write down $Z=b_{C}^{2}C+Z^{\sharp}$, where the effective 1-cycle $Z^{\sharp}$ does not contain the curve $C$ as a component. Obviously,
$\mathop{\rm deg}Z=4b_{C}^{2}+\mathop{\rm deg}Z^{\sharp}=4n^{2}$
and
$\mathop{\rm mult}\nolimits_{o_{1}}Z\geqslant 3b_{C}^{2}+\mathop{\rm deg}Z^{\sharp}=4n^{2}-b_{C}^{2}.$
Combining these estimates, we obtain the inequality
$18n^{2}-24n\theta+10\theta^{2}\leqslant 4n^{2}-\frac{1}{36}(15n-10\theta)^{2},$
which after easy transformations takes the form $460\theta^{2}-1164n\theta+729n^{2}\leqslant 0$, where $\theta\in[n,\frac{9}{8}n]$. However, the quadratic function $460t^{2}-1164t+729$ is decreasing on the interval $1,\frac{9}{8}]$ and its value at the right hand end of this interval is $\frac{108}{64}>0$. This contradiction excludes the case $\theta\leqslant\frac{9}{8}n$.
Assume now that $\theta>\frac{9}{8}n$. In that case $3n-2\theta<\frac{2}{3}\theta$, so that, making the estimates (7) and (8) worse, we set $m_{1}=m_{2}=\frac{2}{3}\theta$. From there it follows that
$b_{C}\geqslant\frac{1}{3}(7\theta-6n)$
and $\mathop{\rm mult}\nolimits_{o_{1}}Z\geqslant\frac{26}{9}\theta^{2}$. Again we write down $Z=b_{C}^{2}C+Z^{\sharp}$ and obtain (in the same way as above) the inequality $139\theta^{2}-300n\theta+162n^{2}\leqslant 0$, where $\theta\geqslant\frac{9}{8}n$. However the quadratic function $139t^{2}-300t+162$ is strictly increasing for $t\geqslant\frac{9}{8}$, and its value for $t=\frac{9}{8}$ is $\frac{27}{64}>0$ (up to a factor, it is the same value as above, since for $\theta=\frac{9}{8}n$ we have $\frac{2}{3}\theta=3n-2\theta$). Contradiction.
Proof of Proposition 2.4 in the case of an irreducible plane quartic is complete.
5.5. The case of a plane cubic and a line. In the case (2) we use the constructions and notations of Subsection 3.2, in certain situations modifying the notations, which will be specially noted. The point $p_{1}\in Q_{2}$ and the infinitely near (to $p_{1}$) point $p_{2}$ have the same meaning as in Subsection 5.2. The curve $C^{1}$ is non-singular. Since the case (2) corresponds to finitely many generators $R$ (and finitely many planes $P$), we can say more about the curve $C$: it is a cubic curve with a nodal singularity $o_{1}$, that is, it has two branches at this point, which meet transversally at the point $o_{1}$. This is true for all generators $R$ in the case (2) for a (Zariski) general quartic of the first type. Therefore, the strict transform $C^{2}$ intersects transversally the exceptional divisor ${\mathbb{E}}_{2}$ and the quadric $Q_{2}$ at the point $p_{3}$, where $p_{3}\neq p_{1}$. Since $(P^{2}\circ V_{2})$ does not contain the exceptional line $P^{2}\cap{\mathbb{E}}_{2}$ (because the equality $\mathop{\rm mult}\nolimits_{o_{2}}(P^{1}\circ V^{1})=2$ holds), we conclude that the line
$p_{1},p_{3}]=P^{2}\cap{\mathbb{E}}_{2},$
containing the points $p_{1},p_{3}$ in ${\mathbb{E}}_{2}\cong{\mathbb{P}}^{3}$, is not contained in the non-singular quadric $Q_{2}$. It follows that $p_{3}\not\in T_{p_{1}}Q_{2}$. Set $b_{C}=\mathop{\rm mult}\nolimits_{C}\Sigma$. Now we can state a stronger version of Proposition 5.1.
Proposition 5.2. The following estimate holds:
$\nu_{2}+\nu_{3}+b_{C}\leqslant 3m_{2}.$ (9)
Proof. We use the notations of Proposition 5.1. Obviously, $\mathop{\rm mult}\nolimits_{p_{3}}Y=b^{*}\geqslant b_{C}$. If $p_{1},p_{2}]\not\subset Q_{2}$, then we consider the section $\Lambda$ of the quadric $Q_{2}$ by a plane, containing the line $p_{1},p_{2}]$ and the point $p_{3}$. This plane is unique precisely because $p_{3}\not\in[p_{1},p_{2}]$. The conic $\Lambda$ is certainly non-singular. Write down
$Y=l\Lambda+Y^{\sharp},$
where the 1-cycle $Y^{\sharp}$ is effective and does not contain $\Lambda$ as a component. We have:
$2(m_{2}-l)=(\Lambda\cdot Y^{\sharp})\geqslant(\Lambda\cdot Y^{\sharp})_{p_{1}}+(\Lambda\cdot Y^{\sharp})_{p_{3}}\geqslant(\nu_{2}-l)+(\nu_{3}-l)+(b_{C}-l),$
so that
$\nu_{2}+\nu_{3}+b_{C}\leqslant 2m_{2}+l\leqslant 3m_{2}.$
Assume that $p_{1},p_{2}]\subset Q_{2}$. Set $L_{1}=[p_{1},p_{2}]$ and denote by the symbol $L_{2}$ the only line in the transversal to $L_{1}$ family of lines on $Q_{2}$, passing through the point $p_{3}$ (that is, if $L_{1}$ is of bi-degree $(1,0)$, then $L_{2}$ is of bi-degree $(0,1)$). Write down
$Y=l_{1}L_{1}+l_{2}L_{2}+Y^{\sharp},$
where the effective 1-cycle $Y^{\sharp}$ does not contain $L_{1}$, $L_{2}$ as components and $l_{1},l_{2}\in{\mathbb{Z}}_{+}$. Obviously, $l_{1}\leqslant m_{1}$ and $l_{2}\leqslant m_{2}$ and $Y^{\sharp}$ is of bi-degree $(m_{2}-l_{1},m_{2}-l_{2})$. Now we get:
$2m_{2}-l_{1}-l_{2}=((L_{1}+L_{2})\cdot Y^{\sharp})\geqslant\mathop{\rm mult}\nolimits_{p_{1}}Y^{\sharp}+\mathop{\rm mult}\nolimits_{p_{2}}(Y^{\sharp})^{(2)}+\mathop{\rm mult}\nolimits_{p_{3}}Y^{\sharp}\geqslant$
$\geqslant\nu_{2}+\nu_{3}+b_{C}-2l_{1}-l_{2},$
whence we obtain
$\nu_{2}+\nu_{3}+b_{C}\leqslant 2m_{2}+l_{1}\leqslant 3m_{2}.$
Q.E.D. for the proposition.
Set $b_{L}=\mathop{\rm mult}\nolimits_{L}\Sigma$. Replacing in the notations of Subsection 3.2 the multiplicity $b$ by $\nu_{2}$, the curve $C_{1}$ by $C$ and the line $C_{2}$ by $L$, we get the effective class
$\xi=nH_{S}-m_{1}E_{1,S}-m_{2}E_{2,S}-\nu_{2}R^{2}_{1}-b_{C}C^{2}-b_{L}L^{2}=$
$(n-b_{C})C^{@}+9n-b_{L}-L^{2}+(2n-m_{1}-\nu_{2})R^{2}_{1}+(n-m_{1})R^{2}_{2}+(2n-m_{1}-m_{2})R_{3}$
on the non-singular surface $S^{2}$. The inequalities $(\xi\cdot C^{2})\geqslant 0$ and $(\xi\cdot L^{2})\geqslant 0$ (see the multiplication table (1)) turn into the estimates
$b_{C}\geqslant\frac{1}{2}(b_{L}+2m_{1}+m_{2}+\nu_{2}-3n),b_{L}\geqslant\frac{1}{2}(b_{C}+m_{1}+\nu_{2}-n).$
Substituting the right hand side of the second inequality for $b_{L}$ in the right hand side of the first one, after easy computations we get
$b_{C}\geqslant\frac{1}{3}(5m_{1}+2m_{2}+3\nu_{2}-7n).$ (10)
Besides, Proposition 5.2 gives the estimates
$m_{1}\geqslant m_{2}\geqslant\frac{1}{3}(\nu_{2}+\nu_{3}+b_{C})$
Substituting $\frac{1}{3}(\nu_{2}+\nu_{3}+b_{C})$ for $m_{1},m_{2}$ in (10), we get
$2b_{C}\geqslant 16\nu_{2}+7\nu_{3}-21n>2n$
(since $\nu_{2}\geqslant\nu_{3}>n$), so that $b_{C}>n$, which contradicts the plane section lemma (Proposition 2.2).
Proof of Proposition 2.4 is complete.
Q.E.D. for Theorems 0.2 and 0.3.
Список литературы
  • [1] Pukhlikov A. V., Birational rigidity of quartic three-folds with a double point of rank 3, Advances in Geometry 25 (2025), No. 4, 447–461.
  • [2] Corti A. and Mella M., Birational geometry of terminal quartic 3-folds. I. Amer. J. Math. 126 (2004), No. 4, 739–761.
  • [3] Ahmadinezhad H. and Kaloghiros A.-S., Non-rigid quartic 3-folds, Compos. Math. 152 (2016), No. 5, 955–983.
  • [4] Pukhlikov A. V., Birational automorphisms of a three-dimensional quartic with an elementary singularity, Math. USSR Sb. 63 (1989), 457–482.
  • [5] Iskovskikh V. A. and Manin Yu. I., Three-dimensional quartics and counterexamples to the Lüroth problem, Math. USSR Sb. 86 (1971), No. 1, 140–166.
  • [6] Mella M., Birational geometry of quartic 3-folds. II, Math. Ann. 330 (2004), No. 1, 107–126.
  • [7] Shramov K. A., Birational automorphisms of nodal quartic threefolds, arXiv:0803.4348, 1–32.
  • [8] Cheltsov I., Factorial threefold hypersurfaces, J. Alg. Geom. 19 (2010), 781–791.
  • [9] Shramov C.A. $\mathbb{Q}$}-factorial three-dimensional quartics, Sbornik: Mathematics 198 (2007), No. 7-8, 1165–1174.
  • [10] Krylov I., Okada T., Paemurru E. and Park J., $2n^{2}$-inequality for $cA_{1}$ points and applications to birational rigidity, Compos. Math. 160 (2024), No. 7, 1551–1595.
  • [11] Akizuki Y. and Nakano S., Note on Kodaira-Spencer’s proof of Lefschetz’s theorem, Proc. Jap. Acad., Ser. A 30 (1954), 266–272.
  • [12] Esnault H. and Viehweg E., Lectures on vanishing theorems, DMV Seminar, Band 20. Birkhäuser Verlag, Basel-Bpstpn-Berlin. 1992. 164 pp.
  • [13] Corti A., Singularities of linear systems and 3-fold birational geometry, in ‘‘Explicit Birational Geometry of Threefolds’’, London Mathematical Society Lecture Note Series 281 (2000), Cambridge University Press, 259–312.
  • [14] Pukhlikov A. V., Essentials of the method of maximal singularities, in ‘‘Explicit Birational Geometry of Threefolds’’, London Math. Soc. Lecture Note Ser. 281 (2000), Cambridge University Press, 73–100.
  • [15] Pukhlikov A., Birationally Rigid Varieties. Mathematical Surveys and Monographs 190, AMS, Providence, RI, 2013, 365 pp.
Department of Mathematical Sciences,
The University of Liverpool
pukh@liverpool.ac.uk

### v2

@@
 The present paper is a follow up of [1].
-We prove factoriality and birational rigidity and describe the group of birational automorphisms of two classes of three-dimensional quartics in ${\mathbb{P}}^{4}$ (the ground field is the field of complex numbers $\mathbb{C}$}:
+We prove factoriality and birational rigidity and describe the group of birational automorphisms of two classes of three-dimensional quartics in ${\mathbb{P}}^{4}$ (the ground field is the field of complex numbers $\mathbb{C}$}):
 (1) the quartics $V\subset{\mathbb{P}}^{4}$ with a unique singular point $o_{1}\in V$, which is a quadratic singularity of rank 3, the blowup $\varphi_{1}\colon V_{1}\to V$ of which yields a variety $V_{1}$ with a unique singularity, a non-degenerate quadratic singularity $o_{2}\in V_{1}$ (so that the blowup $\varphi_{2}\colon V_{2}\to V_{1}$ of the point $o_{2}$ gives a non-singular variety $V_{2}$), — in this paper we will call such quartics the quartics of the first type,
@@
 By Theorem 0.1, there is $n[\Sigma]\geqslant 1$, such that $\Sigma\subset|n[\Sigma]\cdot H|$.
-We say that $\Sigma$ has no maximal singularities, if the pair $\left(V,\frac{1}{n[\Sigma]}D\right)$ is canonical, that is, for every excetional divisor $E$ over $V$ the inequality
+We say that $\Sigma$ has no maximal singularities, if the pair $\left(V,\frac{1}{n[\Sigma]}D\right)$ is canonical, that is, for every exceptional divisor $E$ over $V$ the inequality
 $\mathop{\rm ord}\nolimits_{E}D\leqslant n[\Sigma]\cdot a(E)$
@@
 In contrast to the first two main results, its statement is not the same for the quartics of the first and second type.
-This is because the second singularity of a quartiv of the first type is an infinitely near point and does not generate any additional birational automorphisms.
+This is because the second singularity of a quartic of the first type is an infinitely near point and does not generate any additional birational automorphisms.
 For a quartic $V$ of the first type let $\tau_{0}$ be the Galois involution, generated by the projection $V\dashrightarrow{\mathbb{P}}^{3}$ from the point $o_{1}$ and $\tau_{1}$, …, $\tau_{24}$ are the birational involutions, linked to the 24 lines on $V$, passing through the point $o_{1}$, see [4, 1].
@@
 We know that $L(o_{2})\leqslant 24$.
-Proposition 0.1. For a general quartic of the second type the equality $L(o_{2})\leqslant 24$ holds.
-We denote the subgroup in $\mathop{\rm Bir}V$, generated by these $26+L(o_{2})$ involutions, by the symbol $B(V)$.
+Proposition 0.1. For a general quartic of the second type the equality $L(o_{2})=24$ holds.
+We denote by the symbol $B(V)$ the subgroup in $\mathop{\rm Bir}V$, generated by these $26+L(o_{2})$ involutions.
 Theorem 0.3. The subgroup $B(V)$ is normal in $\mathop{\rm Bir}V$, and there is the exact sequence
@@
 Remark 0.2.
-In the semi-direct product for the group $\mathop{\rm Bir}V$ the finite subgroup $\mathop{\rm Aut}V$ acts on $B(V)$ by conjugation, preserving the point $o_{1}$ and permuting the lines, passing through the point $o_{1}$, for quartics of the first type, and preserving both points $o_{1}$, $o_{2}$ and permuting the lines passing through these points (iside each of the two groups of lines).
+In the semi-direct product for the group $\mathop{\rm Bir}V$ the finite subgroup $\mathop{\rm Aut}V$ acts on $B(V)$ by conjugation, preserving the point $o_{1}$ and permuting the lines, passing through the point $o_{1}$, for quartics of the first type, and preserving both points $o_{1}$, $o_{2}$ and permuting the lines passing through these points (inside each of the two groups of lines).
 Thus in the first family of quartics the subgroup $B(V)$ is freely generated by 25 involutions and in the second family for a general quartic it is freely generated by 50 involutions.
@@
 Irreducible quartic hypersurfaces in ${\mathbb{P}}^{4}$ are parameterized by points of a Zariski open subset in the projective space ${\mathbb{P}}^{69}$, that is, by points of a quasi-projective 69-dimensional variety.
-In [1, Subsection 0.4] the problem of studying the ‘‘birational geography’’ of three-dimensional quartics was formulated: to break the natural parameter space into disjoint quasi-projective subvarieties, consisting of quartics with a similar birational type.
-The natural ‘‘measure of size’’ of these subvarieties is their dimension (or the codimension in ${\mathbb{P}}^{69}$).
+In [1, Subsection 0.4] the problem of studying the “birational geography” of three-dimensional quartics was formulated: to break the natural parameter space into disjoint quasi-projective subvarieties, consisting of quartics with a similar birational type.
+The natural “measure of size” of these subvarieties is their dimension (or the codimension in ${\mathbb{P}}^{69}$).
 The seminal paper [5] that started the theory of birational rigidity described the birational type of non-singular quartics, reducing the subsequent work to investigating the 68-dimensional subset of singular quartics, that is, to the subset of codimension 1.
@@
 It follows from [6] (taking into account [7, 8, 9]) and [1] that the codimension of the complement to the set of factorial birationally rigid three-dimensional quartics is at least 3.
-Non-rigid but ‘‘almost rigid’’ quartics were studies in [2], see also [3].
+Non-rigid but “almost rigid” quartics were studies in [2], see also [3].
 For the families and types of singularities that are already accessible for the modern techniques (however, often very hard to study), see.
@@
 We use the notations of Subsections 0.2, 0.3, considering in detail quartics of the first type.
-Note, in the first place, that the proof of factoriality, presented in [2, Remark 1.2], works for a quartics of the first type $V$, too, as by construction the homogeneous polynomial $q_{4}(z_{*})$ is arbitrary: the section of $V$ by the hypersurface ‘‘at infinity’’ is precisely $\{q_{4}=0\}$.
-Therefore, for a general quartic of the first type its section ‘‘at infinity’’ has Picard number 1.
+Note, in the first place, that the proof of factoriality, presented in [2, Remark 1.2], works for a quartics of the first type $V$, too, as by construction the homogeneous polynomial $q_{4}(z_{*})$ is arbitrary: the section of $V$ by the hypersurface “at infinity” is precisely $\{q_{4}=0\}$.
+Therefore, for a general quartic of the first type its section “at infinity” has Picard number 1.
 However, we give a different proof of factoriality that does not use this fact.
 Note, first of all, that on $X_{2}$ the class of the hypersurface $V_{2}$ is $4H_{\mathbb{P}}-2{\mathbb{E}}_{1}-2{\mathbb{E}}_{2}$.
-This class is nef and big, but not ample, so that the Lefschetz theorem does not apply (in contrast to [1]).
+This class is nef and big, but not ample, so that the Lefschetz theorem does not apply (in contrast to the quartics considered in [1]).
 Furthermore, it is easy to see that all cohomology spaces
@@
 $H^{2}({\mathbb{P}}^{4},{\cal O}_{{\mathbb{P}}^{4}}),\quad H^{2}(X_{1},{\cal O}_{X_{1}}),\quad H^{2}(X_{2},{\cal O}_{X_{2}}),$
-and the six similar spaces for $V$, $V_{1}$, $V_{2}$ are trivial, so that (by the explonential sequences for $X_{2}$ and $V_{2}$) there are the natural isomorphisms
+and the six similar spaces for $V$, $V_{1}$, $V_{2}$ are trivial, so that (by the exponential sequences for $X_{2}$ and $V_{2}$) there are the natural isomorphisms
 $\mathop{\rm Pic}X_{2}\otimes{\mathbb{C}}\cong H^{1}(X_{2},\Omega_{X_{2}})$
@@
 (2.3) the centre of $E$ is a curve $B$ of degree $\geqslant 2$, containing the point $o_{1}$.
-In order to prove Theorem 2.1 we have to investigate all these 8 cases.
+In order to prove Theorem 2.1, we have to investigate all these 8 cases.
 Note that essentially this theorem claims that either (1.1) or (2.1) takes place, and in the latter case only one line is the centre of a maximal singularity.
@@
 The strict transforms on $X_{1}$ of the quadrics in ${\mathbb{P}}^{4}$, cutting out the curve $B$, cut out $B^{1}$ on $X_{1}$.
-Therefore, $W_{2}\cap{\mathbb{E}}_{2}$ is a general plane in ${\mathbb{E}}_{2}\cong{\mathbb{P}}^{3}$, containing the point $B_{2}\cap{\mathbb{E}}_{2}$ (when we add the upper index 2, we mean taking the strict transform on $X_{2}$), so that
+Therefore, $W_{2}\cap{\mathbb{E}}_{2}$ is a general plane in ${\mathbb{E}}_{2}\cong{\mathbb{P}}^{3}$, containing the point $B^{2}\cap{\mathbb{E}}_{2}$ (when we add the upper index 2, we mean taking the strict transform on $X_{2}$), so that
 $W^{2}\cap Q_{2}=S^{2}\cap Q_{2}=(W^{2}\cap{\mathbb{E}}_{2})\cap Q_{2}$
@@
 which implies that $\mu<n$. Q.E.D. for the proposition.
-(In fact, the computations performed above in the singular case and the compuations in [1, Subsection 2.2] in the non-singular case give a stronger estimate $\mu<\frac{1}{4}(-3+\sqrt{41})$, but we do not need it.)
+(In fact, the computations performed above in the singular case and the computations in [1, Subsection 2.2] in the non-singular case give a stronger estimate $\mu<\frac{1}{4}(-3+\sqrt{41})$, but we do not need it.)
 Thus the curve $B$ is contained in some two-dimensional plane $P\ni o_{1}$ in ${\mathbb{P}}^{4}$.
@@
 $\Sigma\subset|nY-m_{1}Q_{1}-m_{2}Q_{2}|$
-is a mobile linear system, however, the inetegers $m_{1}$, $m_{2}$ are not related by the ineqyality $m_{1}\geqslant m_{2}$.
+is a mobile linear system, however, the integers $m_{1}$, $m_{2}$ are not related by the inequality $m_{1}\geqslant m_{2}$.
 Let $L_{1,1},\dots,L_{1,24}$ be the lines on $V$, passing through the point $o_{1}$, and $L_{2,1},\dots,L_{2,L(o_{2})}$ the lines on $V$, passing through the point $o_{2}$.
@@
 which completes the proof of the proposition.
-Proof of Proposition 2.5 is put off until §4, since the point $o_{i}$ and the line $L_{j,l}$ are contained in some plane, which requires the proof of the plane section lemma.
+Proof of Proposition 2.5 we put off until §4, since the point $o_{i}$ and the line $L_{j,l}$ are contained in some plane, which requires the proof of the plane section lemma.
 Proof of Proposition 2.6. Set
@@
 Recall that we consider quartics of the first type, so that $o_{2}$ is the vertex of the cone $Q_{1}$.
-Let $P\ni o_{1}$ be some two-dimensional plane, $P_{1}\subset X_{1}$ its strict transform and $R=P^{1}\cap{\mathbb{E}}_{1}$ the exceptional line of the blow up of the point $o_{1}$ on $P$.
+Let $P\ni o_{1}$ be some two-dimensional plane, $P^{1}\subset X_{1}$ its strict transform and $R=P^{1}\cap{\mathbb{E}}_{1}$ the exceptional line of the blow up of the point $o_{1}$ on $P$.
 If $R$ is not a generator of the cone $Q_{1}$, then $R\not\subset Q_{1}$.
@@
 Its affine part is given by the equation $f(z_{*})|_{P}=0$.
-It is easy to see that $R\subset(P_{1}\circ V^{1})$ if and only if $\mu\geqslant 3$, and $R$ comes into the effective 1-cycle $(P^{1}\circ V^{1})$ with multiplicity $\mu_{2}$.
+It is easy to see that $R\subset(P^{1}\circ V^{1})$ if and only if $\mu\geqslant 3$, and $R$ comes into the effective 1-cycle $(P^{1}\circ V^{1})$ with multiplicity $\mu-2$.
 In order to reduce the number of cases which we have to consider, assume in addition that the quartic $V$ satisfies the following condition of general position: in the notations of Subsection 0.2 the polynomial $q_{3}(z_{*})$ does not vanish identically on all generators of the cone $Q_{1}=\{q_{2}=0\}$ (quartics, for which this condition is violated, form a subset of codimension 2 in the quasi-projective variety of quartics of the first type, and moreover, for every plane $P\ni o_{1}$ the inequality $\mu\leqslant 3$ holds, that is, the effective cycle $(P\circ V)$ is neither $a_{i}L_{i}+a_{j}L_{j}$, where $a_{i}+a_{j}=4$ and $i\neq j$, nor $4L_{i}$ (here $i,j\in\{1,\dots,24\}$).
@@
 Let us list all options for the cycle $(P\circ V)$.
-Each option has its number in the format (Pa.b), where $a=\mu$.
+Each option has its number in the format (a.b), where $a=\mu$.
 We consider the planes, for which $(P\circ V)$ is not an irreducible curve of degree 4, such curves can not be the centre of a maximal singularity, see Subsection 2.2.
@@
 This is all that Proposition 2.2 claims in the case (3.1), so that the plane section lemma is shown for that case.
-Note, that $\theta_{22}=-\frac{3}{24}$, and this agrees with the fact that the line $C_{2}$ can be a maximal curve of the system $\Sigma$.
+Note that $\theta_{22}=-\frac{3}{24}$, and this agrees with the fact that the line $C_{2}$ can be a maximal curve of the system $\Sigma$.
 In order to prove Proposition 2.2 it was sufficient to consider the orthogonal complement to the three classes $R_{1}^{2}$, $R_{2}^{2}$, $R_{3}$.
@@
 Remark 3.1.
-The argument above, when we restrict the linear system $\Sigma$ onto a general surface $S\in|H-P|$ and study the intersection form on the verctor space, the basis of which is formed by the numerical classes of the components of the plane section and the exceptional curves of the resolution of singularities of the surface $S$, is the explicit geometric form of the argument in [4, §6], which was carried out in terms of numerical classes of a certain blow-up of the quartic.
+The argument above, when we restrict the linear system $\Sigma$ onto a general surface $S\in|H-P|$ and study the intersection form on the vector space, the basis of which is formed by the numerical classes of the components of the plane section and the exceptional curves of the resolution of singularities of the surface $S$, is the explicit geometric form of the argument in [4, §6], which was carried out in terms of numerical classes of a certain blow-up of the quartic.
 In [4, §6] it was done in that way in order to avoid considering a large number of possible cases.
-In this paper we restrict the amount of cases by the conditions of general position and so it is easier to prove the plane section lemma for each case, in which the point $o_{2}$ is present, by explicit geometric arguments, resolving singularities of the surface $S$ and analysing the intersection form for a particular set of curves.
-However, the logic of the arguments is the same as in [4, §6]; in particular, the replacement of the classes $C_{1}^{2}$, $C_{2}^{2}$, $R_{1}^{2}$ by the classes $C_{1}^{+}$, $C_{2}^{+}$, $R_{1}^{+}$, orthogonal to the curves $R_{2}^{2}$, $R_{3}$, corresponds to the ‘‘reductions’’ in [4, §6].
+In this paper we restrict the amount of cases by the conditions of general position and so it is easier to prove the plane section lemma for each case, in which the point $o_{2}$ is present, by explicit geometric arguments, resolving singularities of the surface $S$ and studying the intersection form for a particular set of curves.
+However, the logic of the arguments is the same as in [4, §6]; in particular, the replacement of the classes $C_{1}^{2}$, $C_{2}^{2}$, $R_{1}^{2}$ by the classes $C_{1}^{+}$, $C_{2}^{+}$, $R_{1}^{+}$, orthogonal to the curves $R_{2}^{2}$, $R_{3}$, corresponds to the “reductions” in [4, §6].
 For the remaining cases (2.1) and (3.2) we only point out the necessary changes in the arguments for the case (3.1) considered above.
@@
 $H_{S}=C_{1}^{2}+C_{2}^{2}+R_{1}^{2}+R_{2}^{2}+2R_{3}$
-(where in contrast to the previous cases $R_{1}\neq R$: the line $R\subset{\mathbb{E}}_{1}$ intersects the cone $Q_{1}$ at the only point, the vertex $o_{2}$), besides,
+(where, in contrast to the previous cases, $R_{1}\neq R$: the line $R\subset{\mathbb{E}}_{1}$ intersects the cone $Q_{1}$ at the only point, the vertex $o_{2}$), besides,
 $E_{1,S}=R_{1}^{2}+R_{2}^{2}+R_{3}\quad\mbox{and}\quad E_{2,S}=R_{3}.$
@@
 Here $L=C_{3}$, and the multiplication table for the five curves $R_{1}^{2},C_{1}^{2},C_{2}^{2},C_{3}^{2},R_{2}^{2}$ is given by the matrix (we did not give it in Subsection 4.2):
-$\left(\begin{array}[]{ccccc}-2&1&1&0&0\\ 1&-2&1&1&1\\ 1&1&-2&1&0\\ 0&1&1&-2&1\\ 0&1&0&1&-1\end{array}\right).$
+$\left(\begin{array}[]{ccccc}-2&1&1&0&0\\ 1&-2&1&1&1\\ 1&1&-2&1&0\\ 0&1&1&-2&1\\ 0&1&0&1&-2\end{array}\right).$
 The orthogonal complement to $C_{1}^{2}$, $C_{2}^{2}$ is given by the formulas
@@
 From these lemmas we get the equalities $q_{3,0}=-2$, $q_{4,0}=1$. We also assume that the second equality of Lemma 4.2 holds.
-Lemma 4.3 In the coordinates $(z_{1},z_{2},u,z_{4})$ the quadratic part of the polynomial $f$ at the point $o_{2}$ is
+Lemma 4.3. In the coordinates $(z_{1},z_{2},u,z_{4})$ the quadratic part of the polynomial $f$ at the point $o_{2}$ is
 $w_{2}=u^{2}+uq_{4,1}+(z_{1}^{2}+z_{2}^{2}+q_{3,2}+q_{4,2}),$ (4)
@@
 $\sigma_{i}\colon V^{(i)}\to V^{(i-1)}$
-be the sequence of blow-ups of curves $B_{i-1}\subset V^{(i-1)}$ with the exceptional divisors $E^{(i)}\subset V^{(i)}$, $E^{(i)}=\sigma_{i}^{(-1)}(B_{i-1})$, where $i=2,\dots,k+1$, $k\geqslant 2$, such that:
+be the sequence of blow-ups of curves $B_{i-1}\subset V^{(i-1)}$ with the exceptional divisors $E^{(i)}\subset V^{(i)}$, $E^{(i)}=\sigma_{i}^{-1}(B_{i-1})$, where $i=2,\dots,k+1$, $k\geqslant 2$, such that:
 $V^{(1)}=V_{2}$ and $B_{1}=R^{2}$,
@@
 The Noether-Fano inequality takes the simplified form
-$m_{1}+\nu_{2}+\dots\nu_{k+1}>(k+1)n.$
+$m_{1}+\nu_{2}+\dots+\nu_{k+1}>(k+1)n.$
 Let $Z=(D_{1}\circ D_{2})$ be the self-intersection of the system $\Sigma$, where $D_{1},D_{2}\in\Sigma$ are general divisors.
-This is and effective 1-cycle on $V$, for which, according to the technique of counting multiplicities, see [15, Chapter 2, Proposition 2.4, Corollary 2.5], and also [15, Chapter 2, Subsection 2.2], [1, §3], we have the estimate
+This is and effective 1-cycle on $V$, for which, according to the technique of counting multiplicities, see [15, Chapter 2, Proposition 2.4, Corollary 2.5], and the computation at the end of [15, Chapter 2, Subsection 2.2], and also [1, §3], we have the estimate
 $\mathop{\rm mult}\nolimits_{o_{1}}Z>\frac{2(k+1)^{2}}{2k+1}n^{2}.$
@@
 $\left(\sigma_{2}|_{Q_{2}^{(2)}}\right)^{*}D^{2}|_{Q_{2}}-\nu_{2}N.$
-This effective 1-cycle on $Q_{2}^{(2)}$ has multiplicity $\geqslant\nu_{2}$ at the point $p_{2}=B_{2}\cap Q_{2}^{(2)}=B_{2}\cap N$ (the inetrsections are transversal, since $B_{2}$ is a section of the ruled surface $E^{(2)}/B_{1}$).
+This effective 1-cycle on $Q_{2}^{(2)}$ has multiplicity $\geqslant\nu_{3}$ at the point $p_{2}=B_{2}\cap Q_{2}^{(2)}=B_{2}\cap N$ (the intersections are transversal, since $B_{2}$ is a section of the ruled surface $E^{(2)}/B_{1}$).
 Therefore, we obtain the following geometric situation.
@@
 $\nu_{2}+\nu_{3}\leqslant\mathop{\rm mult}\nolimits_{p_{1}}Y+\mathop{\rm mult}\nolimits_{p_{2}}Y^{(2)}=\mathop{\rm mult}\nolimits_{p_{1}}Y^{+}+\mathop{\rm mult}\nolimits_{p_{2}}(Y^{+})^{(2)}+2a\leqslant 3m_{2},$
-where $Y^{(2)}$ means the strict transform. This completes the proof of Proposition 5.1.
+where $Y^{(2)}$ means the corresponding strict transform. This completes the proof of Proposition 5.1.
 5.3. The two techniques of the proof. For a quartic of the first type of general position there are only two options, see Subsections 3.1 and 3.5:
-(1) the case of general position, when $(V\circ P)=V\cap P$ is an irreducible curve $C$ of degree 4 with the triple point $o_{1}$, so that $o_{2}\in C^{1}$ is a non-singular point (this case takes place for allmost all generators $R$, except for finitely many of them),
+(1) the case of general position, when $(V\circ P)=V\cap P$ is an irreducible curve $C$ of degree 4 with the triple point $o_{1}$, so that $o_{2}\in C^{1}$ is a non-singular point (this case takes place for almost all generators $R$, except for finitely many of them),
 (2) $(V\circ P)=C+L$, where $C$ is an irreducible cubic curve with the double point $o_{1}$, $L\ni o_{1}$ is a line, and moreover, $o_{2}\not\in L^{1}$ and $o_{2}\in C^{1}$ is a non-singular point.
 We will exclude these two cases in different ways: the method of excluding the case (1) does not work in the case (2), and the method of excluding the case (2) does not apply to finitely many generators $R$ and planes $P$, for which the curve $C$ has at the point $o_{1}$ two branches, a non-singular one and a cuspidal one, and moreover, the point $o_{2}$ lies exactly on the cuspidal branch.
-The case (1) will be excluded by means of the ‘‘quadratic’’ technique, which is more traditional and simpler.
+The case (1) will be excluded by means of the “quadratic” technique, which is more traditional and simpler.
 Our approach to the proof of Proposition 2.4 in the case (1) is to estimate the input of the curve $C$ to the self-intersection $Z$ of the system $\Sigma$; we will show that this input is sufficiently big, which will make it possible to obtain a contradiction by means of the standard technique of counting multiplicities:
@@
 $b_{C}\geqslant\frac{1}{2}(3m_{1}+m_{2}+2\nu_{2}-4n).$ (8)
-(from the inequalities (5) and (6) it follows that the right hand side is strictly positive).
-It is easy to check that for a fixed value $\nu_{2}+\nu_{3}=2\theta$ the estimates (7) and (8) are the worst for $\nu_{2}=\nu_{3}=\theta$ (recall that $\nu_{2}\geqslant\nu_{3}$: on a non-singular variety multiplicities do not increase), so that we may assume that this equality takes place.
+(It follows from the inequalities (5) and (6) that the right hand side is strictly positive.) It is easy to check that for a fixed value $\nu_{2}+\nu_{3}=2\theta$ the estimates (7) and (8) are the worst for $\nu_{2}=\nu_{3}=\theta$ (recall that $\nu_{2}\geqslant\nu_{3}$: on a non-singular variety multiplicities do not increase), so that we may assume that this equality takes place.
 Consider first the case $\theta\in(n,\frac{9}{8}n]$. In that case the inequality
@@
 $\mathop{\rm mult}\nolimits_{o_{1}}Z\geqslant 18n^{2}-24n\theta+10\theta^{2}.$
-Write down $Z=b_{C}^{2}C+Z^{\sharp}$, where the effective 1-cycle $Z^{\sharp}$ does not contain the curve $C$ as a component. Obviously,
-$\mathop{\rm deg}Z=4b_{C}^{2}+\mathop{\rm deg}Z^{\sharp}=4n^{2}$
+Write down $Z=\beta C+Z^{\sharp}$, where the effective 1-cycle $Z^{\sharp}$ does not contain the curve $C$ as a component and $\beta\geqslant b_{C}^{2}$. Obviously,
+$\mathop{\rm deg}Z=4\beta+\mathop{\rm deg}Z^{\sharp}=4n^{2}$
 and
-$\mathop{\rm mult}\nolimits_{o_{1}}Z\geqslant 3b_{C}^{2}+\mathop{\rm deg}Z^{\sharp}=4n^{2}-b_{C}^{2}.$
+$\mathop{\rm mult}\nolimits_{o_{1}}Z\leqslant 3\beta+\mathop{\rm deg}Z^{\sharp}=4n^{2}-\beta.$
 Combining these estimates, we obtain the inequality
@@
 and $\mathop{\rm mult}\nolimits_{o_{1}}Z\geqslant\frac{26}{9}\theta^{2}$.
-Again we write down $Z=b_{C}^{2}C+Z^{\sharp}$ and obtain (in the same way as above) the inequality $139\theta^{2}-300n\theta+162n^{2}\leqslant 0$, where $\theta\geqslant\frac{9}{8}n$.
+Again we write down $Z=\beta C+Z^{\sharp}$ with $\beta\geqslant b_{C}^{2}$ and obtain (in the same way as above) the inequality $139\theta^{2}-300n\theta+162n^{2}\leqslant 0$, where $\theta\geqslant\frac{9}{8}n$.
 However the quadratic function $139t^{2}-300t+162$ is strictly increasing for $t\geqslant\frac{9}{8}$, and its value for $t=\frac{9}{8}$ is $\frac{27}{64}>0$ (up to a factor, it is the same value as above, since for $\theta=\frac{9}{8}n$ we have $\frac{2}{3}\theta=3n-2\theta$).
@@
 where the effective 1-cycle $Y^{\sharp}$ does not contain $L_{1}$, $L_{2}$ as components and $l_{1},l_{2}\in{\mathbb{Z}}_{+}$.
-Obviously, $l_{1}\leqslant m_{1}$ and $l_{2}\leqslant m_{2}$ and $Y^{\sharp}$ is of bi-degree $(m_{2}-l_{1},m_{2}-l_{2})$.
+Obviously, $l_{1}\leqslant m_{2}$ and $l_{2}\leqslant m_{2}$ and $Y^{\sharp}$ is of bi-degree $(m_{2}-l_{1},m_{2}-l_{2})$.
 Now we get:
@@
 $\xi=nH_{S}-m_{1}E_{1,S}-m_{2}E_{2,S}-\nu_{2}R^{2}_{1}-b_{C}C^{2}-b_{L}L^{2}=$
-$(n-b_{C})C^{@}+9n-b_{L}-L^{2}+(2n-m_{1}-\nu_{2})R^{2}_{1}+(n-m_{1})R^{2}_{2}+(2n-m_{1}-m_{2})R_{3}$
-on the non-singular surface $S^{2}$. The inequalities $(\xi\cdot C^{2})\geqslant 0$ and $(\xi\cdot L^{2})\geqslant 0$ (see the multiplication table (1)) turn into the estimates
-$b_{C}\geqslant\frac{1}{2}(b_{L}+2m_{1}+m_{2}+\nu_{2}-3n),b_{L}\geqslant\frac{1}{2}(b_{C}+m_{1}+\nu_{2}-n).$
+$(n-b_{C})C^{2}+(n-b_{L})L^{2}+(2n-m_{1}-\nu_{2})R^{2}_{1}+(n-m_{1})R^{2}_{2}+(2n-m_{1}-m_{2})R_{3}$
+on the non-singular surface $S^{2}$. The inequalities $(\xi\cdot C^{2})\geqslant 0$ and $(\xi\cdot L^{2})\geqslant 0$ turn into the estimates (see the multiplication table (1))
+$b_{C}\geqslant\frac{1}{2}(b_{L}+2m_{1}+m_{2}+\nu_{2}-3n),\quad b_{L}\geqslant\frac{1}{2}(b_{C}+m_{1}+\nu_{2}-n).$
 Substituting the right hand side of the second inequality for $b_{L}$ in the right hand side of the first one, after easy computations we get
@@
 Q.E.D. for Theorems 0.2 and 0.3.
-Список литературы
-  • [1] Pukhlikov A. V., Birational rigidity of quartic three-folds with a double point of rank 3, Advances in Geometry 25 (2025), No. 4, 447–461.
-  • [2] Corti A. and Mella M., Birational geometry of terminal quartic 3-folds. I. Amer. J. Math. 126 (2004), No. 4, 739–761.
-  • [3] Ahmadinezhad H. and Kaloghiros A.-S., Non-rigid quartic 3-folds, Compos. Math. 152 (2016), No. 5, 955–983.
-  • [4] Pukhlikov A. V., Birational automorphisms of a three-dimensional quartic with an elementary singularity, Math. USSR Sb. 63 (1989), 457–482.
-  • [5] Iskovskikh V. A. and Manin Yu. I., Three-dimensional quartics and counterexamples to the Lüroth problem, Math. USSR Sb. 86 (1971), No. 1, 140–166.
-  • [6] Mella M., Birational geometry of quartic 3-folds. II, Math. Ann. 330 (2004), No. 1, 107–126.
-  • [7] Shramov K. A., Birational automorphisms of nodal quartic threefolds, arXiv:0803.4348, 1–32.
-  • [8] Cheltsov I., Factorial threefold hypersurfaces, J. Alg. Geom. 19 (2010), 781–791.
-  • [9] Shramov C.A. $\mathbb{Q}$}-factorial three-dimensional quartics, Sbornik: Mathematics 198 (2007), No. 7-8, 1165–1174.
-  • [10] Krylov I., Okada T., Paemurru E. and Park J., $2n^{2}$-inequality for $cA_{1}$ points and applications to birational rigidity, Compos. Math. 160 (2024), No. 7, 1551–1595.
-  • [11] Akizuki Y. and Nakano S., Note on Kodaira-Spencer’s proof of Lefschetz’s theorem, Proc. Jap. Acad., Ser. A 30 (1954), 266–272.
-  • [12] Esnault H. and Viehweg E., Lectures on vanishing theorems, DMV Seminar, Band 20. Birkhäuser Verlag, Basel-Bpstpn-Berlin. 1992. 164 pp.
-  • [13] Corti A., Singularities of linear systems and 3-fold birational geometry, in ‘‘Explicit Birational Geometry of Threefolds’’, London Mathematical Society Lecture Note Series 281 (2000), Cambridge University Press, 259–312.
-  • [14] Pukhlikov A.
-V., Essentials of the method of maximal singularities, in ‘‘Explicit Birational Geometry of Threefolds’’, London Math.
-Soc.
-Lecture Note Ser.
-281 (2000), Cambridge University Press, 73–100.
-  • [15] Pukhlikov A., Birationally Rigid Varieties. Mathematical Surveys and Monographs 190, AMS, Providence, RI, 2013, 365 pp.
-Department of Mathematical Sciences,
-The University of Liverpool
-pukh@liverpool.ac.uk